A Bibliography of Publications in The Microsoft Systems Journal

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

FAX: +1 801 585 1640, +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@ieee.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 November 2002 Version 1.10

Title word cross-reference

/ [Ano97a].

16-bit [Eds94c, Fin94, Fin95c, Gel93, One93].

2 [Chi90, Dun90a, FWB90, FL90, Hil90, OH90, Sha90a, Sha90d, Yao91c]. 2.0 [Yao98]. 21ST [RC98].

3 [Pro95a]. 3-D [Pro95a]. 32 [Ric93e]. 32-Bit [Yao91d, Cus92, Dun90a, One93, Pet92b, Pie96b, Ric94a, Sch92b, Sch93a, She91]. 3D [Cla98a, Fos96, Fos98].

4.0 [Ano96b, Arm96a, Arm96b, CE97, Rob98c].

5.0 [BBM98, EE97a, EE97b, One97a, One97c, Pie97b, Pie97c, Qui98, RC98, Rob97b, She97].

6.0 [DiL98g, Mas98].

80386- [Dun90a]. 80486-based [Dun90a]. 80x86 [Yao90].

95 [Eds95b, One95c, Pro95d]. 97 [Kir97c, Zim97b]. 98 [One97a, One97c].

Accept [Ric92a]. Access [Ano96g, MTM96, Pro94b, Rau97, Tru93b]. Accessing [Sha90a]. Accurately [Jon91]. Active [Box97a, Box97b, DDBM98, Ano96a, Box96a, CE97, Rau96b]. ActiveX [Ano97a, Box97c, Box97d, Box97e, Box98a, Box98b, Box98c, EE97a, EE97b, Pla97b, Tru96f].

ActiveX-based [Pla97b]. ActiveX/ [Box97d, Box98a, Box98c]. ActiveX/COM [Box97c, Box97e, Box98b]. Activity [Chi90]. Adapt [Hal91a]. Adapting [Bil92, Gre90, Hal91b]. Add [Ric95b, Zim97b]. Add-ins [Zim97b]. Added [Thi94]. Adding [Adl90a, Kle92a, Pet90a]. Additional [Adl90f]. Address [Ric94a]. Addressability [Sha90b]. Advance [Kin95]. Advanced [Cla98a, Pro95a, Tru95b, Yao91d]. Advantage [Cun98, Wei97]. Age [Pet92b]. Agree [Pla97a]. Aid [Wei96, Wei97]. All-Encompassing [Jac97]. Allocation [Yao91a, Yao91b]. Almost [Adl90f, DiL94c, FWB90]. Analyzing [EE98, Sha91a]. Animation [Ano93, Eds95a]. Annihilating [Rob97c]. Answer [Roz96]. Any [Chi91, Fin95c, Rau96a, Tru93c, Yao98, Kir97b]. Anything [Sch92b]. Anywhere [Sch92b]. Apartment [Pla97b]. API [AMT93, Ano96g, MR93, Pet90d, Ric95b, Roz96, Ume98, Yao91d]. APls [Fin98a]. App [EE95, Fin95b, Pla97a, Ric95b]. Application [Adl90a, Adl90f, Adl91e, Ano96i, Azi94, Cha98b, Chi92, Dix91a, Kel98, One95c, PS91, Pet90d, Pro95g, Ric92a, Ric96a, Rob97c, Tru94, Yao91a, Yao91b]. Applications [BBES98, BD98, DiL98f, Dun90b, Eds92b, Eds94c, Eis90, EMR91, Fin95a, FWB90, Gel93, Hel95, Mas98, One93, Pie96b, Ric96b, Ric98a, Sha90a, Sha90d, Sha92a, She91, SC97, Smi92, Tru94, Wel90e, Wel90f, Wel92]. Approach [Hol95]. Apps [Ano96b, Bla96, Cla98a, Kir97c, Nic96, Not96, Not97, Pat98, RTT97, Zim97a]. Appwizards [One97b]. Architect [Wil93]. Architecting [SCP97]. Architecture [Ano95c, Dun90a, Ker90, Kin95, Pro96a]. Architectures [Yao90]. Areas [Yao92a]. Art [Bol92]. ASP [Pat98]. Attention [One97b]. Audio [Pet93].

Auto [Ric93a]. Auto-Reset [Ric93a].

Automated [Nor92]. Automatic [Pro93a]. Automatically [One96]. Automation [Bro95]. aware [PS92].

Back [Roz96]. Bag [DiL98g, Eds95b].

Bars [Ano95a, Pro95c].

Based [EE95, Pat98, Sha90b, SCP97, Adl90a, Adl90c, Bol98, BBES98, Coo96, Dun90a, Eds92b, Pie93c, Pla97b, Sha92a, Ano96e, Fin95a, Dun90b]. Basic [Ano96b, EE97a, EE97b, Not96, Not97, Pat98, Rob97b, Rob97d, OH90, Ano96i, EE96, EMR91, Tru93b, Tru93c, Tru94, Tru95b, Tru96e]. Be [Cla98b]. Beasties [Eds93]. Behind [One97b, SCP97]. Best [Cap93]. Better [Ano95e, BBES98, Chi94]. Between [Bro93c, Pie93d, Ric92c]. Beyond [Eds94b, Hol96c]. Bidirectional [Gre90]. Bit [Yao91d, Cus92, Dun90a, Eds94c, Fin94, Fin95c, Gel93, One93, Pet92b, Pie96b, Ric94a, Sch92b, Sch93a, She91]. Bitmaps [Adl91c, Mir91]. Bits [Pet92a, Pie95a]. Block [Cus95]. Bona [Sch93a]. Boost [Pla97b]. Borland [Sha91b, Tru95a]. Bottlenecks [One94]. Boxes [Adl91a, Azi94, Pro95f, Woo90]. Bridges [Bro93c]. Brings [DiL98g, Pet93, Qui98, Rau96b]. Browser [Rob98c]. Browsing [Rau96b]. Brushes [Pro95d]. Bugs [Eds92b, Rob97c]. Bugslayer [Ano98a, Rob97a, Rob97c, Rob98a, Rob98b]. Build [McK96, Pat98]. Building [Ano96e, Box97a, BBES98, BD98, Fai94a, Fin95a, Huc96, Kir97c, Lan95, Not96, Not97, One95b, SC97, Wel90a]. Builds [Ano95e, Chi94]. Built [Mey93a]. Built-in [Mey93a]. Bull [Kle92b]. Button [Adl91e]. Buttons [DiL98f].

C [Ano95e, Chi93, Chi94, EE95, Hol93, Lan95, MS91, Mit96, Nic96, Pro95g, Sha92b, She97, Thi94, Ano98b, Ber90, Bro93d, Com90, DeL98, DiL93a, DiL93b, DiL94a, DiL94b, DiL95b, DiL95g, DiL95e, DiL95a, DiL95d, DiL95h, DiL95c, DiL95f, DiL96j, DiL96e, DiL96f, DiL96g, DiL96h, DiL96i, DiL96a, DiL96b, DiL96d, DiL97d, DiL97e, DiL97f, DiL97g, DiL97h, DiL97a, DiL98a, DiL98c, DiL98d, DiL98e, DiL98g, Dil96c, Dil97b, Dil97c, Dil98b, Goo92, Mas98, MTM96, Mey95, One97b, Sha91a, Sha91b, Sha92a]. C/C [Sha92b, DiL93a, DiL94a, DiL94b, DiL95b, DiL95g, DiL95e, DiL95a, DiL95d, DiL95h, DiL95c, DiL95f, DiL96e, DiL96f, DiL96i, DiL96j, DiL96g, DiL96h]. Cache [Cus95]. Cake [Bla96]. Call [MS91, Pie93c, Sch93b]. Calls [Fin95a]. Can [SC97]. Capabilities [EE96, Sha93, Tru93b, Yao91d]. Careful [Yao91a]. Case [Pet92a]. CE [Ano97e, FR97, Yao98, Bol98, Jon98]. CE-based [Bol98]. Century [RC98]. Certain [Kim98]. Changes [Cha98a]. Channel [CE97]. Checkerboard [Pet90c]. Checkers [Pet90c, Pet90b, Pet90a]. Chicago [Ano95b, Ano95a, Ano95c, Eds94a, Eds94c, Kin94, Kin95, One95a, Pie95a]. Chief [Wil93]. Choice [Cha98b]. Clandestine [Fin98b, Fin98c]. Class [DiL93b, Eds96, Fai94b, Fos96, Huc96, Sha92c, Sha93]. Classes [Ano96f, Bla96, Chi95, Pro95g, Mey93a, Sha92b, Sha92c]. Cleanup [Yao91a]. Clearer [Goo95]. Cleverly [Ano95b, Eds94a]. Clien [EE96]. Clien/Server [EE96]. Client [Dix91a, Dix91b, Fin95a, Fin95b, Ker90, Mit96, Not96, Not97, PS91, Smi92, Ano96i]. Client-Server [Dix91a, Dix91b, Ker90, PS91, Smi92]. Client/ [Not96, Not97]. Client/Server [Fin95a, Fin95b, Ano96i]. Clipboard [Wel90a]. Code [Ano96h, Fin94, Goo92, Hol93, Jam90a, MS91, One93, Pie93c, Pro97a, Pro96c, Pro96d, Pro97b, Pro97c, Pro97d, Sha91a, Sha92a].

Codewright [Ric92b].

Coding [Ano95b, Eds94a]. Com [Box96a, Box97d, Box98a, Box98c, Ano96a, Box96b, Box97c, Box97e, Box97a, Box97b, Box98b, BBES98, Box98d, Bro96a, Bro96b, Cha98a, Kir97b, Kir97a, Rau96a, She97, Sko98]. COM- [BBES98]. Combining [Sha90b]. Comes [Cus92, Pet92b]. Command [DiL95i]. Commands [Jam90a]. Common [Ano96b, Fin98b, Fin98c, Arm96a, Arm96b]. Communication [Gre90]. Compact [Box97a]. Comparing [Adl91e]. Compiled [OH90]. Compiler [Hol93]. Compilers [Sha91a]. Complete [Sha91c]. Complex [BBM98, Ker90, Pro94b]. Component [Bro93b, Bro96a, Bro96b, Kir96, Lan95, Ano96d]. Components [Kir97b, Pat98, Sko98]. Comprehensive [Goo92, Goo95, Sha91b]. Compression [Pro93a]. Computer [Roz96]. Computing [Jon98]. Configuration [Pro93a]. Configure [One96]. Configuring [Ano96g]. Conform [One95b]. Considerations [Bro95]. Containers [Chi95, DiL94c]. Contending [Pro95e]. Control [Adl90e, Ano95a, EE97a, EE97b, Eds95a, Jam90a, Kel98, Key92, Lan95, Pro94b, Ric92d, Ric98b, Tru93a, Fin98b, Fin98c]. Controller [Bro95]. Controls [Adl91a, Ano96c, Arm96a, Arm96b, Bro92, McK95, McK96, Pro95f, Tru93c, Wel90d, Ano96b, Fin98b, Fin98c]. Converting [Mir91]. Cool [DiL98f, Wei97]. Cooperative [Hal91b]. Coordinate [Ric93a]. Corpulent [Pie93a]. Corral [Adl91e]. Course [AS95, Pie97a]. CPUMETER [Chi90]. Crash [AS95, Pie97a]. Create [EMR91, Zim97a]. Creating [DiL94c, Dix91a, Dix91b, Fos96, Ric93b, Wel90b, Wel90c, Wel92, Woo90]. Creation [EE97a, EE97b].

Critical [Ric93e]. Cross [EE95]. Cross-Development [EE95]. Curtain [One97b]. Curve [Nic96]. Custom [Bro92, Cun98, Key92, Lan95, Ric98a, Wel90d, Yao91b]. Customization [BP98].

D [Pro95a]. DAO [MTM96]. Data [Ano96j, Bro93a, Chi92, EE98, Hol96a, Hol96b, Hol97, MTM96, Rau97, Ric93d, Wel90b, Wel90c]. Data-Entry [Wel90b, Wel90c]. Database [Rau96a, Tru93b]. Databases [MTM96]. DB [Ano96d, Rau96a]. DCOM [EE98]. DCs [Pro95d]. DDE [Bro93c, Pet91b, Sha90c, Sha90d, Smi92]. Debugger [Rob97d]. Debuggers [Mir93]. Debugging [Chi91, Hil90, Mey93a, Mey93b, Sch97]. Decompression [Ric92d]. Default [Chi92]. Definition [Hlu98]. Delivers [Chi95]. Delphi [Tru95a]. Demand [Sch92a]. Demand-Paged [Sch92a]. Deployment [Pat98]. Deposits [Pie96b]. depth [Mef90]. Depths [Pie97a]. Design [Ano96a, Ano96f, Bro95, Bro96a, Bro96b, Cus95, Hol95, Ric97a, Sko98]. Designers [Pla97a]. Designing [Bro92, Cap93, PS91]. Desktop [Ano96i, Fai98, Rau96b, Ric96a, CE97]. Destroying [Ric93b]. Details [Cus95]. Developer [Ano97b, Gea90, Hlu98, Lan95, Zim97b]. Developers [DiL98g, Yao91d]. Developing [Fin95b, Hel95, Mas98, Nor92, PS92, She97]. Development [Adl91d, Ber90, EE95, Hil90, Ker90, Kir97a, Pet90d, Qui98, Sha90e, Sha91b, Thi91b]. Device [Adl91b, Huc96, Mir91, One95b, Ric96c, Sch92b, Smi91]. Device-Independent [Mir91]. Devices [Ano96g, One96, Pro94b, Smi91]. Diagnostic [Mey93a, Mey93b]. Dialog [Adl91a, Azi94, Pro95f].

DIBs [DiL97k, DiL97i, DiL97j].

Difference [Pie93d]. Digital [Pet93]. Direct [Ano96e, Eds95a, Eds96]. Directinput [Eds96, Cla98b]. Directory [DDBM98]. DirectSound [Cla98a]. DirectX [Cla98a, Cla98b, Fos98]. Discomfort [EE97a, EE97b]. Disk [Pro93a]. Disparate [Yao90]. Dispensers [Cun98]. Distributed [Box96b, Cha98b, Cun98, Kir97c, RTT97]. DLL [Arm96a, Arm96b, Chi95, Fin95c, Pie97b, Pie97c, Ric94a]. DLLs [Eis90, Fai94a, Key92, Ric92d]. Do [Kim98, Sch92b, Sha92a]. Document [Ano96a, Pet90d, Tru96f, Pro96a]. Document/View [Pro96a]. DOS-based [Dun90b]. Down [Sha92a]. Download [Kir96, Ano96d]. Dr. [Tru93a]. Drag [Ric92a]. Drag-and-Drop [Ric92a]. Dramatically [Nic96]. Draw [Fos98]. Drawing [Pet90c]. Drive [Fin94]. Driver [One97a, One97c]. Drivers [Fin98a, Huc96, One95b, Sch92b, Smi91]. Drop [Ric92a]. Dynamic [Azi94, SC97, Wel90a]. Dynamic-Link [Wel90a]. Dynamically [One95c].

Easier [Pet91b, Sha91c]. Easily [Sch93b, Zim97a]. Easy [DiL94c, Kir97b, Tru95b]. Edit [Adl90b]. Edition [EE95, EE96, Not96, Not97]. Editor [Ric92b]. Editors [Ano95e, Chi94]. Effective [BBES98, Cap93]. Efficient [Cha98b]. Either [Bro93d]. Embedded [Smi91]. Embedding [Kle92a, Kle92b]. Encompassing [Jac97]. Engine [MTM96]. Enhance [Azi94, Tru93c]. Enhanced [Pet91a, Sch92a]. Enhancing [Adl90b]. Enterprise [Ano96b, EE96, Not96, Not97]. Enters [Chi93].

Entry [Tru95a, Tru95b, Wel90b, Wel90c]. Environment [Ber90, Dun90b, EMR91, Sha91b, Sha91c, Sha90e]. Error [Goo95]. Events [Jon91, Ric93a, Rob98c]. Everything [Ric92a]. Examination [Goo92]. Examining [Jam90b]. Exception [Goo95, Mey95, Pie97a]. Exceptions [Ano95e, Chi94]. Excess [Pie93a]. Exciting [Jon98]. Exclusive [Huc96, Pie95b, Ric95b]. Executable [Pie94a, Wel91]. Executables [Pie93a]. Experience [Mit96]. Exploit [Bol92, Cam97, Ric97a]. Exploiting [Dun90a, One96, Sko98]. Exploration [Mef90]. Explore [Yao92a]. Explorer [Arm96a, CE97, Rob98c, Cam96]. Exploring [Adl91b, Cla98a, EE96, Pet92b, Sch92a]. Exposes [Fin98b, Fin98c]. Extend [Key92, One95c, Ric96a, Zim97b]. Extended [Hal91b]. Extending [Cam96, Eds94b, Eis90, Wel90d]. Extends [Fai98]. Extensible [Wel90a, Zim97a]. Extension [DDBM98]. Extensions [Ano96j, Cam96, Pet91a, Rau96b, Ano96d, Mir96]. Exterminate [Eds92b]. Extra [Tru93c]. Eye [Rob98c].

Face [Cap93]. Facilities [Mey93a, Mey93b, Ric97a, Sha90a, Sha90d]. Fall [Pla97a]. Far [Sha90b]. Farewell [Pet91b]. Fashionable [Pla97a]. Fast [Cha98b, Fos98, McK96]. Faster [Ano95e, Chi94]. Fat [Pie93a, One95a]. Fatty [Pie96b]. Features [Ano95e, Box96b, Chi94, Mas98, Pie97b, Pie97c, Qui98, Tru95b]. Feedback [Cla98b]. Fide [Sch93a]. Fields [Adl90b]. File [One95a, Pie94a, Ric93d, RC98, Wel91]. Filenames [One95a]. Files [Ric92a, Ric93d, Ric95b, Thi92].

Filling [Thi91a]. Filter [Wel92]. Filters [BP98]. Fingers [Tru96e]. First [Wei96, Wei97]. Fonts [Adl91c]. Footprint [Bol98]. Force [Cla98b]. Foreign [Mir91]. Formal [Bil92]. Format [Pie94a]. Formats [Wel91]. Formatted [Adl90b]. Forward [Cap93]. Foundation [Mey93a, Sha92b, Sha92c, Sha93]. Four [Mir93]. Framework [Pro95g]. Free [Pla97a]. Frontier [Hol96c]. FTP [AS95]. Full [DiL98g, Pet93]. Full-Motion [Pet93]. Fun [DiL93b, DiL97k, DiL97i, DiL97j]. Function [Fin95c]. Functionality [Key92]. Functions [MS91, Pie93b, Pie95b, Sch93b, Thi92, Tru96e]. Fundamental [Ano93, Com90]. Fundamentals [Yao92b]. Fusing [Ric96b].

Gadgets [Ano95b, Ano95a, Eds94a]. Gain [Kel98]. Game [Eds95a, Pet90a]. Game-Playing [Pet90a]. Games [Eds96]. Gamut [DiL97k, DiL97i, DiL97j]. Gates [Pie93c]. GDI [Pet92b]. Generation [Kin95]. Generators [Adl91e]. Generic [Fin94, Sch93b]. Genius [Ric92c]. Get [DiL93b, Eds96, Fos98, McK95]. Getting [AS95]. Give [DiL98f, Pla97b]. Gives [Cla98a, Eds95a, Yao91d]. Global [Pie93b]. GMMF [Ric95b]. Go [Hol96c, Sch92b]. Going [Eds94b]. Goodies [DiL97k, DiL97i, DiL97j, Eds95b]. Gopher [AS95]. Gotchas [Eds95b]. GPI [Pet90c]. Grab [Eds95b]. Grand [Cus92]. Graphical [EMR91, Sha91c]. Graphics [Adl91b, Pet92b, Pro95a, Wel92]. Grappling [Cla98b]. Great [Thi91a]. Growable [Ric95b]. Guide [Hlu98, Jac97].

Hair [Gel93].

Handheld [FR97, Ric96c, Yao98].

Handling [Goo95, Jam90a, Mey95, Pie97a]. Handy [Chi91]. Hardware [Eds95a].

Harnessing [Tru93b]. Heap [Pie93b, Ric93c]. HeapPeep [Chi92]. Hello [Pet91b]. Help [Adl90a, DiL93b, Fai98]. helps [RTT97]. Hide [Hol96a, Hol96b, Hol97]. Hiding [Ano96j]. Hints [Eds92a]. Home [DiL98g]. Hood [Ano98c, Pie93b, Pie95c, Pie95d, Pie95e, Pie95f, Pie96a, Pie96c, Pie96d, Pie96e, Pie96f, Pie96g, Pie96h, Pie96i, Pie96j, Pie96l, Pie96m, Pie97e, Pie97f, Pie97g, Pie97h, Pie97i, Pie97j, Pie97k, Pie97l, Pie98a, Pie98b, Pie98c, Pie98d, Pie98e, Pie98f, Pie96k, Pie97d]. Hook [Fin95c]. Horns [Kle92b]. Hot [DiL98f, Pla97a]. House [Box98d]. HTML [Fai98]. HTTP [Ano96e, Coo96, Sko98]. HTTP-Based [Ano96e]. Hybrid [Ano95c]. Hygiene [Yao91a]. Hypertext [Adl90a]. Hypertext-based [Adl90a].

I/O [MS91]. IDEs [Sch97]. IE [Arm96b, Ano96b]. II [Adl90d, Ano95a, Ano96a, Arm96a, Bro93d, Bro96a, Cla98a, DiL97i, EE97b, Fin98c, Hal94b, Hol96b, Kle92a, Mey93b, Not97, Pet90c, Pie97c, Pro95d, Rob97d, Wel90b]. III [Adl90e, Ano95d, Bro93a, DiL97j, Hol97, Pet90b]. IIS [BP98]. Image [Ano95b, Eds94a]. Impaired [Wei96, Wei97]. Implement [Rob97b]. Implementation [Cus95, Hol93]. Implementations [Hol96c]. Implementing [Bro93a, Bro93d, Bro95, DDBM98, One97a, Sha90c]. Import [Wel92]. Improve [Yao91b]. Improved [Arm96b, Chi95]. Improves [Yao91a]. In-depth [Mef90]. Includes [Ano97b]. Independent [Mir91]. Industrial [Ric92b]. Industrial-Strength [Ric92b]. Ingenious [Rob97b]. INJLIB [Ric94a]. Input [Ano95d, Ano96e, Ric92c].

Ins [Zim97a, Zim97b]. Installable [Wel90d]. Installer [Kel98]. Integrated [Ber90, Pro93a]. Integrating [Adl91a]. Integration [Ano96c]. Interactive [Bla96, McK96, Sha90e]. Interface [Adl91b, Ano95a, Azi94, Cap93, DiL98f, Eds94a, Hlu98, Pat92, Pet90d, Pro94b, Pro95a, Rau96a, Wel90d, Adl90b]. Interfaces [Bak93, Hol96a, Hol96b, Hol97, Rau97]. Interfacing [OH90]. Interleaved [Pet93]. Internal [Kin94]. International [FL90]. Internationalization [Hal91a, Hal94a, Hal94b]. Internet [Arm96a, AS95, Ano96c, Ano96d, CE97, Kir96, McK96, Mit96, Nic96, Rau96b, RTT97, Rob98c, SCP97]. Internet-Based [SCP97]. Interoperability [Tru94]. Interpreters [BP98]. Interrupt [Mef90]. Interview [Wil93]. Introducing [Box96b, Bro93b, Kir97c, One95a, Pro95a, Rob97c, Ano96d]. Introduction [Gea90, Pat92, Ric93c, Yao92b]. Investigating [Ano95c, Kle92a, Kle92b, Sch97]. ISAPI [Ano96f, Bla96, BP98]. Isolate [Eds92b]. IV [Adl91a, Pet90a, Pro95e].

Jet [MTM96]. Joy [Box97a, Eds96].

Just [Tru93c, Yao98].

Keeping [Rob98c].

Kernel [Fin98a, Huc96, Sha90a, Sha90d]. Kernel-Mode [Huc96]. Key [Ric92c]. Keyboard [Mef90, Pro95e, Ric92c]. Killer [Hel95]. Kit [FWB90, Lan95]. Knockout [Chi93]. Know [Ric96c].

LAN [Dix91a, Dix91b, PS91].

Language [FL90, Hlu98, Hol93, Kir97b]. Languages [Hal94b]. Last [Sch93a]. Last-Write [Sch93a]. Layer [Fin94].

Layout [BBM98]. Learn [Chi90, Mir96]. Learned [Tru93a]. Learning [Adl90c, Adl90d, Adl90e, Adl91a, Adl91b, Adl91c, Nic96]. Less [Sha92a, Thi91a]. Lets [Chi91]. Level [BP98]. Leveraging [Mit96]. Libraries [DiL93b, Wel90a, Wel92]. Library [Box97a, Fai94b, Huc96, MS91, Sha93, She91, Smi92, Wel90b, Wel90c, Yao92a]. Life [Fin98b, Fin98c]. Light [Chi92]. Line [Ume98]. Link [Wel90a]. Linking [Kle92a, Kle92b]. Links [Ano97b]. Liposuction [Pie93a, Pie96b]. List [Eds92a, Woo90]. Lists [Ano95b, Eds94a]. Load [Ric94a]. Loaded [One95c]. Loading [Ano97a]. Locales [Hal94b]. Look [DiL98f, Kin95, Mas98, Pie93b]. Love [Tru93a].

Mac [EE95]. Machine [Chi90]. Macros [Eis90, Zim97b]. Made [DiL94c, Kir97a]. Maintain [Pie93b]. Maintenance [Kel98]. Make [Hol96a, Hol96b, Hol97, Ric92a]. Makes [Box97a, Kir97b]. Man [One97b]. Manage [Rau97]. Management [Kin94, Pie95a, Ric93c, Smi92, Yao90, Yao92b]. Manager [Adl90b, Cus95, Adl90g, Dix91a, Dix91b, Hil90, OH90, PS91, Pet90b, Sha90a, Sha90c]. Managing [Ric93b, Thi91b]. Manipulate [Ric98b]. Manipulation [Ric93d]. Manual [Ric93a]. Manual-Reset [Ric93a]. Mapped [Ric93d, Ric95b]. May [Cla98b]. Maze [DiL95i]. MDI [Adl90e, Pet90d]. Meandering [DiL95i]. Measure [Chi90]. Media [Pro94b]. Memory [Adl90f, Bol98, Kin94, Pie95a, Pro93a, Ric93c, Ric93d, Ric95b, Sch92a, Yao90, Yao91b, Yao92b]. Memory-Mapped [Ric93d, Ric95b]. Memphis [Cam97]. Menu [DiL98f]. Menuing [Adl90d]. Menus [Pro95c].

Message [Adl90c, Cha98b, DiL95i]. Message-based [Adl90c]. Messages [Chi91]. Messaging [Ano95c]. Methodologies [Wel90f]. Methodology [Zim97a]. MFC [Ano95d, Ano96b, Ano96c, Ano96e, Ano96f, Ano96j, Bla96, Chi95, Coo96, DiL93b, DiL94c, DiL95i, DiL97k, DiL97i, DiL97j, DiL98g, Fai94a, Fos96, Hol95, Hol96c, Mey93b, Pro95e, Pro95d, Pro95c, Pro95f, Pro95g, Pro96a, Pro96b, Thi94, Wei96]. MicroQuill’s [Yao91c]. Microsoft [Adl90a, Ano97b, Arm96a, Arm96b, Ber90, Cha98a, Cha98b, EMR91, Gea90, Goo92, Hal91b, Mey93a, Nor92, RTT97, Ric96c, Roz96, Sha92b, Sha92c, Sha93, Yao98]. Microsoft’s [Wil93]. Migrating [Eds93]. Migration [FWB90]. Minimizing [Bol98]. Misbehave [Kim98]. Mix [One93]. Mobile [Jon98]. Mode [Dun90b, Fin98a, Huc96, One95a, Pet91b, Sch92a]. Model [Bro93b, Cha98a, Dix91b, Kir97b, One97a, One97c, Pla97a, Pla97b]. Models [Chi95]. Modem [Mir96]. Modify [SC97]. Monitor [Cam97, Fin95c]. Monitoring [Ric98a, Rob98c]. More-Sophisticated [Mey93b]. Most [DiL93b]. Motion [Pet93]. Mouse [Ano95d]. Moving [Pet90b]. MS [Bol92, Pro93a, Pro93b, Pro94a, Pro95b, Thi91a, Thi92]. MS-DOS [Bol92, Pro93a, Pro93b, Pro94a, Pro95b, Thi91a, Thi92]. MSJ [Wil93]. MSMQ [BD98]. MSN [Ano97e, SCP97]. MTS [BD98, Cun98, Pat98]. Multilanguage [BBM98, Sch97]. Multimedia [Fin95b, Pet91a, Pro94b]. Multipage [Wel90c]. Multiple [Cam97, Pet90d, Rob97b, Rob97d, Wei96]. Multiprocessing [Cus92]. Multitasking [Hal91b]. Multithreading [Wei97]. Multitiered [Not96, Not97]. Mutexes [Ric93e]. Myriad [Rau97].

Name [Cam96, Ano96d]. Named [Dix91a]. Namespace [DDBM98]. National [FL90]. Navigate [Mey95]. Near [Rob97c, Sha90b]. Need [FR97, Ric96c]. NetBIOS [PS92, Pat92]. Network [AMT93, Ano97b, Bak93, Bro93c, Dix91b, EE98, PS92, Pat92]. Network-aware [PS92]. Networking [Jon98]. Next [BP98, Kin95]. No [Adl90f, One97b]. Noise [Eds96]. NT [BBM98, Fin94, One97a, One97c, Pie96a, Pie97b, Pie97c, Qui98, RC98, Ano96f, Ano96j, Bak93, Cam97, Cus92, Cus95, Eds93, Fin95a, Fin98a, Gel93, Hal94a, Hal94b, Hel95, Huc96, Kim98, One94, Pet92b, Pro95a, Ric93b, Ric93d, Ric97a, Ric98a, Ric98b, Yao92b]. NT-Based [Fin95a]. NTM [Box96b].

O [MS91]. Object [Bro93b, Bro93d, Hol95, Jam90b, Kir97a, Kle92a, Kle92b, Sha90c, Wel90f]. Object-Oriented [Hol95, Jam90b, Wel90f]. Objects [Ano96g, Bro93b, Bro93a, Hol96a, Hol96b, Hol97, MTM96, She97, SC97, Tru96f, Ano96a, Box97a]. OCX [Chi95]. Off [Eis90]. Off-the-Shelf [Eis90]. Offers [Jon98, Pro93a]. OK [Adl91e]. OLE [Ano96d, Ano96c, Ano96f, Box95d, Box95a, Box95c, Box95b, Box96b, Box96c, Box96d, Box96e, Box96f, Bro93b, Bro93a, Bro93d, Bro95, Bro96b, DiL94c, Lan95, McK95, McK96, Pet91b, Rau96a, Tru93b, Tru94, Wil93]. Online [Fai98, SCP97, Ano97b]. Open [Fin98a]. OpenGL [Ano96b, Fos96, Pro95a]. Operating [Ric97a, Yao91d]. Optimization [Sha91a, Wel90e]. Oriented [Hol95, Jam90b, Kir97a, Wel90f]. OS/2 [Chi90, Dun90a, FWB90, FL90, Hil90, OH90, Sha90a, Sha90d, Yao91c]. Our [Fin95c, Huc96, Pie96b, Ric95b, Sch93b, Zim97a, Jac97]. Overview [Adl91d, Bak93, Ric92b].

OWL [Fai94a, Fai94b].

Own [Ano96a, Hol96a, Hol96b, Hol96c, Hol97, One97b, Zim97b].

Paces [Fai94b, Mir93, Tru96f]. Packets [EE98]. Paged [Sch92a]. Pages [Pla97b]. Pain [EE97a, EE97b, Fos96]. Painlessly [FWB90]. Palettes [DiL97k, DiL97i, DiL97j]. Palt [Pro95c]. Panel [Key92]. Pantheon [Tru95a]. Paradigm [Adl90c]. Parallel [Gre90]. Pare [Sha92a]. Part [Adl90c, Adl90d, Adl90e, Adl91a, Adl91b, Adl91c, Ano95b, Ano95a, Ano95d, Ano96a, Ano96c, Ano96e, Arm96a, Bro93b, Bro93a, Bro93d, Bro96a, Cla98a, DiL97i, DiL97j, EE97b, Eds94a, Fin98b, Fin98c, Hal94a, Hal94b, Hol96b, Hol97, Kle92a, Kle92b, Mey93b, Not97, Pet90c, Pet90b, Pet90a, Pie97b, Pie97c, Pro95a, Pro95e, Pro95d, Pro95f, Pro96a, Pro96b, Rob97b, Rob97d, Wel90b]. Pay [One97b]. PC [FR97, Mef90, MR93, One95a, Pet91a, Pet93, Thi91b]. PCs [Bro93c, Yao98]. PE [Pie94a]. Peering [Pie94a]. Pegasus [Ric96c]. Pens [Pro95d]. Performance [One94, Ric98a, Yao91c]. Performance-tuning [One94]. Persistence [Ano96a, Hol96c]. Perspective [Pie97b, Pie97c]. Piece [Bla96]. Pieces [Pet90c, Pet90b]. Pipes [Dix91a]. Play [Ano96g, One96]. Playing [Pet90a]. Plug [AMT93, Ano96g, One96]. PM [Pet90a, Wel90e]. Pointer [Sha90b]. Pointers [Sha90b]. Poking [Pie96a]. Pop [Fin98a]. Port [EE95, FWB90, Gel93, Gre90]. Portable [Cus92, Pie94a]. Porting [Eds93, She91]. Portion [PS91]. Possibilities [Jon98]. Post [Chi91]. Postman [Chi91]. Potpourri [Thi94]. Power [Tru93c]. Preparing [Eds94c].

Presentation [Hil90, Adl90b, Adl90g, OH90, Pet90b, Sha90a, Sha90c]. Press [Mir96, Ume98]. Preview [CE97, Sha92b]. Previewing [Arm96a, Arm96b, Pro96b, RC98]. Previously [Yao92a]. Primer [Ano96a, Ano96c, Ano96e, Ano96f, Bro92, Fai94a]. Primitive [Fos98]. Print [Pro96b]. Printing [Adl91c, Pro96b]. Privileged [Fin98a, Pie93c]. Problems [Bro96a, Bro96b]. Procedure [Fin95a]. Procedures [Nor92]. Process [Jam90a, Pie95a, Ric94a, She97]. Processes [Hal91b, Ric93b]. Processing [Ano95d, Goo95]. ProcHook [Fin95c]. Program [Bol98, FR97, Hal91a, Hol95, One94, Pie93c, Ric92d, Ric98b]. Programmer [Ano97c, Ano97d, Ano98d, Jac97, Pie96a, Pie97b, Pie97c, SW97a, SW97b, SW97c, SW98a, SW98b, Thi91a, Thi94, Tru96e, Tru96f, Tru96a, Tru96b, Tru96c, Tru96d, Tru96g]. Programming [AMT93, Ano95d, Bak93, BBES98, Cha98a, Com90, Eds92a, Eds95b, Hal94a, Kir97b, Pat92, Pro95e, Pro95d, Pro95c, Pro95f, Pro96a, Pro96b, Sha91c, Sha92c, Tru95a]. Programs [Adl90f, Bil92, Bro93c, FL90, PS92, Ric92c, Sch93a, Yao91c]. Property [Pro95f]. Protected [Dun90b, One95a]. Protected-Mode [Dun90b]. Protocol [EE98]. Prototypes [Tru93c]. Prototyping [Adl90g]. Provides [Ber90, Sha91b, Tru95b]. Providing [Sha90d]. Puts [Tru96f]. Putting [Cap93, Fai94b, Mir93].

Q [Box95a, Box95c, Box95b, Box96c, Box96d, Box96e, Box96f, DiL93a, DiL94a, DiL94b, DiL95b, DiL95g, DiL95e, DiL95a, DiL95d, DiL95h, DiL95c, DiL95f, DiL96j, DiL96e, DiL96f, DiL96g, DiL96h, DiL96i, DiL96a, DiL96b, DiL96d, DiL97d, DiL97e, DiL97f, DiL97g, Dil96c, Dil97b, Dil97c, Fin93, Pie94b, Pie94c, Pie95g, Pie95h, Pie95i, Pie95j, Pro93b, Pro94a, Pro95b, Ric94b, Ric95c, Ric95d, Ric96g, Ric96h, Ric97c, Ric97d, Box95d, Box96a, Ric95a, Ric96d, Ric96e, Ric96i, Ric96f]. Quality [Hol93]. Queue [Cha98b]. QuickC [Sha91c]. QuickPascal [Jam90b].

R [Pie96a, Mir96, Ric96i, Ric96f]. Raid [Mir93]. Reach [MR93]. Ready [McK95]. Real [EMR91, Pet91b]. Recommendations [Com90]. Reduces [Nic96]. Refined [McK95]. Reliable [Ano97a]. Relieves [EE97a, EE97b]. Remote [Fin95a]. Remove [Pie93a, Pie96b]. Removing [One94]. Rendering [Fos98]. Reporting [Wel90b, Wel90c]. Reprogramming [Jon91]. Requires [Ric92c]. Reset [Ric93a]. Resource [Cun98, Wel91, Yao91a]. Resources [Adl90d, Hal94b, Yao91b]. Responsible [Hol96a, Hol96b, Hol97]. Reusable [DiL97k, Fos96, Sko98]. Revealed [Wel91]. Rewriting [Hol95]. RichEdit [Ano95a]. Ring [Chi93]. Ripping [Gel93]. Riptoe [Jac97]. Robust [McK95]. Roll [Ano96a, Hol96c]. Routines [Mef90, Ric93c]. Routing [DiL95i]. RPCs [Fin95b]. Run [Pie93c, Sch93a]. Runtime [Kir97a, SC97].

Safe [Kir96]. Safer [Thi92].

Scalable [RTT97]. Scenes [SCP97]. Scores [Chi93]. Scribble [Hol95]. Script [BP98]. Scripts [BBM98, Nor92]. Scrolling [Woo90]. SDK [Ano97e, Eds95a, FR97]. Sections [Ric93e]. Segment [Chi92]. Segmentor [Yao91c]. Semaphores [Ric93e]. Serious [AMT93].

Server [Ano96e, Ano97a, Coo96, Dix91a, Dix91b, EE96, Fai94a, Fin95a, Fin95b, Hel95, Ker90, Kle92a, Not96, Not97, PS91, Smi92, Ano96i, Cha98a, RTT97]. Server-Style [Fai94a]. Servers [DiL94c]. Service [Dix91b, Kir96, Mef90, Ric97a, Ric98b, SCP97]. Services [DDBM98, FL90, Kim98, Kir97a, One96, Ric98b, Sch93b]. Set [Fin98a]. Setup [Kel98, Ric92d]. Seven [BBES98]. Seventeen [Eds94c]. Sharing [Ric93d]. Shed [Chi92]. Sheets [Pro95f]. Shelf [Eis90]. Shell [Ano96c, Ric96a, Ric96b]. Shepherd [Ano97d]. Ship [McK95]. Shoot [Adl91e]. Shoot-out [Adl91e]. Signal [Jam90a]. Simple [Bro93d, Bro95, Coo96, Fos98, Kir97a]. Simplifies [Pet90d, She97]. Simplify [Azi94, BD98, Mey93a, Pro94b, Ric93d, Smi91, Smi92]. Simplifying [Ker90]. Simulating [Ric92c]. Size [Sha90b]. Slay [Eds93]. Small [Sha90b]. Smart [McK96, Zim97a]. Smell [Pro95g]. Snap [Zim97a]. Snap- [Zim97a]. Sock [Ano96i]. Sockets [AMT93, Coo96]. Software [Bro96a, Bro96b, FWB90, Kir97a, Lan95, Thi91b]. Solve [Bro96b]. Solves [Bro96a]. Some [Chi92]. Someone [MR93]. Sophisticated [Mey93b]. Sound [Ano96e, Cla98a, Pet91a]. Source [Ric92a]. Sources [Rau97]. Space [Cam96, Ric94a, Ano96d]. Spawn [Adl90f]. Special [Ric97a]. Speech [Ano96g, Roz96]. Speeds [Huc96]. Speedy [Eds95a]. Sprite [Ano93]. Spy [Fin98b, Fin98c]. Standard [MS91]. Startup [Goo92]. State [Bol92]. State-of-the-Art [Bol92]. Status [Ano95a, Pro95c]. Stay [Ume98]. Stepping [Pie95a]. stocked [Kir97c]. Stop [Tru93a]. Story [Ano97e]. Strangerelease [Tru93a]. Strategies [Bol92]. Strategy [Pet90a]. Streamline [Ric92d]. Strength [Ric92b].

Stroke [Ric92c].

Structured [Goo95, Pie97a]. Studio [Kir97c, Zim97b]. Style [Com90, Fai94a]. Subclassing [DiL97k, DiL97i, DiL97j]. Subsegment [Yao91b]. Subsystem [Kin94]. Support [Adl90e, Cam97, FL90, Hal91a, Kle92a, Smi91, Tru93b, Woo90]. Supporting [BBM98, Dun90b, Yao90]. Surfing [Kir96]. Survey [Adl90g, Thi91a]. Surveying [One97c, Sha93]. Survival [Hlu98]. Swinging [Chi93]. Synchronization [Kin94]. Synchronizing [Ric93e]. System [Adl90d, Cun98, Fai98, One95a, Pie97b, Pie97c, Ric96b, Ric97a, RC98, Yao91d, Adl90a]. Systems [Dun90a, Tru93a].

Take [BP98, Cun98, Wei97].

Taking [Kle92b]. Talk [Rau96a, Roz96]. TAPI [Qui98]. Tasks [Pie93d]. Tastes [Thi91a]. Technical [Chi93]. Techniques [Ano93, Bil92, Bol92, Eds92a, Eds94c, Jam90b, Mey93b, Sha91a, Wel90e]. Technologies [Kim98]. Technology [Mit96]. Telephony [MR93, Qui98, Ume98]. Template [Box97a]. Templates [Ano95e, Chi94]. Ten [Eds92a, Eds93]. Test [Fin94, Nor92]. TestBasic [Nor92]. Testing [Bil92, Nor92]. Text [BBM98]. Their [Hol96b, Hol97, Mir93, Tru96f, Hol96a]. There [Wel91]. Thin [Mit96]. Thoughts [Thi91b]. Thread [Pie95a, Wei96, Wei97]. Thread-Impaired [Wei96]. Threading [Pla97a, Pla97b]. Threads [Chi90, Ric93b, Ric93a, Ric93e, Rob97b, Rob97d, Wei96]. Thunk [Fin94, One93]. Timer [Jon91]. Timing [Jon91]. Tips [BBES98, DiL93b, Eds93, Hel95]. TM [Tru96f]. Tony [Wil93].

Tool [Chi91, Hil90]. Toolbars [Ano95a, Ano96i, Pro95c, Ric96a]. ToolBook [Sha90e]. Toolbox [Kir97c]. ToolHelp [Yao92a]. TOOLHELP32 [Pie95b]. Tools [Adl90g, Adl91d, FR97, One94, Pie96b, Thi94, Tru95a]. ToolTip [Jac97]. Tooltips [Ano97e, Jac97]. Top [Eds92a, Eds93]. Total [Eds96]. Touch [MR93, Sha92c]. Tour [Cus92, Pie94a, Pie95b]. TRACER [Hil90]. Track [Eds92b]. Transaction [Cha98a, RTT97]. Transactional [BD98]. Transfer [Bro93a]. Transform [Sha91a]. Treacherous [Mey95]. Treats [DiL98g]. TreeViews [Ano95b, Eds94a]. Tricks [DiL98g, Eds92a, Hel95]. TrueType [Pet91b]. TSRs [Bol92]. Tuning [Wel90e, Yao91c, One94].

Uncharted [Yao92a]. Unconstrained [One95a]. Unconventional [Thi91b]. Understanding [EE98, Hlu98]. Unicode [Hal94a]. Unified [Rau96b]. Uniform [Bro93a]. Unimodem [Ume98]. Universal [One93, Rau97]. Unloading [Ano97a]. Unusual [Smi91]. Upcoming [Ric96c]. Use [BD98, Chi90, Hal91a, Ker90, Sha91a, Wel90a, Yao91b]. User [Adl90b, Azi94, Cap93, Hol96a, Hol96b, Hol97]. Users [SC97]. Using [Adl90a, Adl90f, Bro93d, Coo96, DDBM98, DiL94c, Dix91b, EMR91, Fin94, Fin95a, Fin95b, FL90, Hol95, Jam90b, Mas98, MS91, MTM96, PS92, Pat92, Pet90c, Pie93c, Pie96b, Pro95g, Rau96a, Ric93a, Ric93e, Ric94a, Sch93a, Sch93b, Sha92a, She97, Smi92, Wei96, Wel90f, Yao92a].

V [Adl91b, Pro95c]. Value [Thi94]. Value-Added [Thi94]. Version [Ber90, Dun90a, Gea90, Goo92, Pet91b, Ric92d, Sha93, Thi91a, Tru93a, Wel90c]. VI [Adl91c, Pro95f]. Video [Fin95b, Pet91a, Pet93]. View [DDBM98, Fos96, Pie96a, Pro96a]. Viewer [Wel90a]. Viewpoint [Gea90]. VII [Ano96e, Pro96a]. VIII [Ano96c, Pro96b]. Virtual [Cus95, Ric93c, Sch92a, Sch92b]. Virtualized [Woo90]. Visual [Ano97d, Ano98d, DiL98g, EMR91, Mas98, Pat98, Rob97d, SW97a, SW97b, SW97c, SW98a, SW98b, Tru95a, Tru96e, Tru96f, Tru96a, Tru96b, Tru96c, Tru96d, Tru96g, Ano97c, Ano95e, Ano96b, Ano96i, Chi93, Chi94, EE95, EE96, EE97a, EE97b, Hol93, Kir97c, Lan95, Mit96, Nic96, Not96, Not97, Pro95g, Rob97b, She97, Thi94, Tru93b, Tru93c, Tru94, Tru95b]. VMM [Sch93b]. Voice [Ano96j, Mir96]. VxD [One96, Sch93b]. VxDs [One95c].

Wake [Pro95g]. Waters [Mey95]. Way [Rau96a]. Ways [Rob97b, Wei97]. Web [Fai98, AS95, Bla96, DDBM98, Kir96, Pla97b]. Well [Kir97c]. Wicked [Ano96h, Pro97a, Pro96c, Pro96d, Pro97b, Pro97c, Pro97d]. Win [Ric96d, Ric97b, Ric98c, Ric95a, Ano96i, Ric93e, Ric95c, Ric96e, Ric96f, Ric97c, Ric97d]. Win32 [Fin94, Goo95, Kim98, Kin94, One97a, One97c, Pie94a, Pie97a, Ric93a, Ric93c, Ric94b, Ric95d, Ric96g, Ric96i, Ric96h, Ric97e, Ric98d, Ric98e, Rob97d]. Win32s [One93, Sch93a]. Window [Chi91]. Windowing [Ano95c]. Windows [Jon98, Yao98, Adl90c, Adl90d, Adl90e, Adl90g, Adl91a, Adl91b, Adl91c, Adl91d, AMT93, Ano93, Ano95d, Ano96c, Ano96f, Ano96j, Ano97e, Bak93, Bil92, BBM98, Bol98, Box96b, Bro93b, Bro93c, Bro93d, Cam96, Cam97, Chi91, Chi92, Coo96, Cus92, Cus95, DiL93b, EE95, Eds92a, Eds92b, Eds93, Eds94b, Eds95b, Eds95a, Eis90, EMR91, Fai94a, Fai94b, Fin93, Fin94, Fin95a, Fin95b, Fin95c, Fin98a, Fin98b, Fin98c, FR97, FWB90, Gea90, Gel93, Hal91b, Hal91a, Hal94a, Hal94b, Hel95, Huc96, Jon91, Kel98, Ker90, Key92, Kim98, Kin95, MS91, McK95, Mir91, Mir93, MR93, Mir96, One94, One95b, One95c, One97a, One97c, PS92, Pet91a, Pet91b, Pet92b, Pet93, Pie93b, Pie93c, Pie93d, Pie94b, Pie94c, Pie95b, Pie95g, Pie95h, Pie95i, Pie95j, Pie96a, Pie97b, Pie97c, Pro95a]. Windows [Pro95e, Pro95d, Pro95c, Pro95f, Pro96a, Pro96b, Qui98, Ric92b, Ric93b, Ric93d, Ric96a, Ric96b, Ric97a, Ric98a, RC98, Ric98b, Sch92a, Sch92b, Sch93a, Sha90c, Sha90e, Sha91b, Sha91c, Sha92a, Sha92b, Sha92c, She91, Smi91, Tru95b, Wel90b, Wel90c, Wel90d, Wel90e, Wel90f, Wel91, Wel92, Woo90, Yao90, Yao91a, Yao91b, Yao91c, Yao91d, Yao92a, Yao92b]. Windows-Based [EE95, Eds92b, Pie93c, Sha92a]. Wingo [Ano97d]. WINIO [MS91]. Winlet [Sko98]. WINMEM32 [She91]. WinSock [Tru96e]. Wire [EE98]. Wired [AS95]. Within [Adl90f, Sha90a]. Without [Fos96, Gel93, Jon91]. Wizards [Zim97b]. WLLs [Eds94b]. Word [Eds94b]. WordBasic [Eds94b]. Workgroups [Bro93c]. Working [Pro95d, Thi92]. World [Eds96]. World-Class [Eds96]. Worldwide [Hal91a]. Worrying [Tru93a]. WOW [Fin94]. Write [Coo96, FL90, Kir97b, One97b, RTT97, Sch93a]. Writing [Bla96, Bol92, BP98, Nic96, Ric98b, Rob97d].

X [Ano96a, Box96a, Box97b, Rau96b]. X/Com [Box96a, Box97b].

References

Adler:1990:AHH

[Adl90a] Marc Adler. Adding hypertextbased help to your application using the Microsoft Help System. Microsoft Systems Journal, 5(3):15–??, May 1990. CODEN MSJOED. ISSN 0889-9932.

Adler:1990:EPM

[Adl90b] Marc Adler. Enhancing the Presentation Manager User Interface with formatted edit fields. Microsoft Systems Journal, 5(1): 35–??, January 1990. CODEN MSJOED. ISSN 0889-9932.

Adler:1990:LWTa

[Adl90c] Marc Adler. Learning Windows part I: The message-based paradigm. Microsoft Systems Journal, 5(4):85–??, July 1990. CODEN MSJOED. ISSN 0889-9932.

Adler:1990:LWTb

[Adl90d] Marc Adler. Learning Windows part II: Resources and the menuing system. Microsoft Systems Journal, 5(5):75–??, September 1990. CODEN MSJOED. ISSN 0889-9932.

Adler:1990:LWTc

[Adl90e] Marc Adler. Learning Windows part III: Control Windows and MDI support. Microsoft Systems Journal, 5(6):67–??, November 1990. CODEN MSJOED. ISSN 0889-9932.

Adler:1990:SPW

[Adl90f] Marc Adler. Spawn programs from within your DOS application using almost no additional memory. Microsoft Systems Journal, 5 (5):33–??, September 1990. CO-DEN MSJOED. ISSN 0889-9932.

Adler:1990:SWT

[Adl90g] Marc Adler. A survey of Windows and Presentation Manager prototyping tools. Microsoft Systems Journal, 5(1):57–??, January 1990. CODEN MSJOED. ISSN 0889-9932.

Adler:1991:LWTa

[Adl91a] Marc Adler. Learning Windows part IV: Integrating controls and dialog boxes. Microsoft Systems Journal, 6(1):97–??, January 1991. CODEN MSJOED. ISSN 0889-9932.

Adler:1991:LWTb

[Adl91b] Marc Adler. Learning Windows part V: Exploring the graphics device interface. Microsoft Systems Journal, 6(2):75–??, March 1991. CODEN MSJOED. ISSN 0889-9932.

Adler:1991:LWTc

[Adl91c] Marc Adler. Learning Windows part VI: Bitmaps, fonts, and printing. Microsoft Systems Journal, 6(3):115–??, May 1991. CO-DEN MSJOED. ISSN 0889-9932.

Adler:1991:OWT

[Adl91d] Marc Adler. An overview of Windows 3.0 development tools. Microsoft Systems Journal, 6(4):99– ??, July 1991. CODEN MSJOED. ISSN 0889-9932.

Adler:1991:SAO

[Adl91e] Marc Adler. Shoot-out at the OK button corral: Comparing application generators. Microsoft Systems Journal, 6(6):87–??, November 1991. CODEN MSJOED. ISSN 0889-9932.

Allard:1993:PSN

[AMT93] J. Allard, Keith Moore, and David Treadwell. Plug into serious network programming with the Windows sockets API. Microsoft Systems Journal, 8(7):35–??, July 1993. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1993:FTS

[Ano93] Anonymous. Fundamental techniques for sprite animation in Windows. Microsoft Systems Journal, 8(12):??, December 1993. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1995:CIG

[Ano95a] Anonymous. Chicago’s interface gadgets, part II: Toolbars, status bars, and the RichEdit control. Microsoft Systems Journal, 10(4):17–??, April 1995. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1995:CCC

[Ano95b] Anonymous. Cleverly coding with Chicago’s gadgets, part I: Image lists and TreeViews. Microsoft Systems Journal, 10(4): 1–??, April 1995. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1995:IHW

[Ano95c] Anonymous. Investigating the hybrid windowing and messaging architecture of Chicago. Microsoft Systems Journal, 10(5):13– ??, May 1995. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1995:PWM

[Ano95d] Anonymous. Programming Windows 95 with MFC, part III: Processing mouse input. Microsoft Systems Journal, 10(8):57–??, August 1995. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1995:VCF

[Ano95e] Anonymous. Visual C++ 2.0 features better editors, faster builds, and C++ exceptions and templates. Microsoft Systems Journal, 10(4):36–??, April 1995. CO-DEN MSJOED. ISSN 0889-9932.

Anonymous:1996:CDP

[Ano96a] Anonymous. COM design primer part II — roll your own persistence — Active X document Ob-Jects. Microsoft Systems Journal, 11(6):??, June 1996. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1996:ICC

[Ano96b] Anonymous. IE 4.0 Common Controls. OpenGL with MFC, enterprise apps with Visual Basic 4.0. Microsoft Systems Journal, 11 (11):??, November 1996. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1996:IOC

[Ano96c] Anonymous. Internet OLE controls. MFC primer part VIII. Windows 95 shell integration. Microsoft Systems Journal, 11(4):??, April 1996. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1996:IOD

[Ano96d] Anonymous. Introducing OLE DB. Name space extensions. Internet component download. Microsoft Systems Journal, 11(7):??, July 1996. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1996:MPP

[Ano96e] Anonymous. MFC primer part VII. building an HTTP-Based server. direct sound and direct input. Microsoft Systems Journal, 11(2):??, February 1996. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1996:OWN

[Ano96f] Anonymous. OLE in Windows NT 4.0. ISAPI classes in MFC 4.1. OLE design primer. Microsoft Systems Journal, 11(5):??, May 1996. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1996:SAA

[Ano96g] Anonymous. Speech API. dat access objects. configuring plug and play devices. Microsoft Systems Journal, 11(1):??, January 1996. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1996:WC

[Ano96h] Anonymous. Wicked code. Microsoft Systems Journal, 11(10): 95–??, October 1996. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1996:WSV

[Ano96i] Anonymous. Win Sock from Visual Basic — application desktop toolbars — Visual Basic client/ server. Microsoft Systems Journal, 11(3):??, March 1996. CO-DEN MSJOED. ISSN 0889-9932.

Anonymous:1996:WNV

[Ano96j] Anonymous. Windows NT 4.0.

voice extensions for Windows. data hiding in MFC. Microsoft Systems Journal, 11(8):??, August 1996. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1997:ARS

[Ano97a] Anonymous. ActiveX — reliable server loading / unloading. Microsoft Systems Journal, 12(1): 93–??, January 1997. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1997:NIL

[Ano97b] Anonymous. Now includes links to Microsoft Developer Network Online! Microsoft Systems Journal, 12(11):??, November 1997. CO-DEN MSJOED. ISSN 0889-9932.

Anonymous:1997:VP

[Ano97c] Anonymous. The visual programmer. Microsoft Systems Journal, 12(2):93–??, February 1997. CO-DEN MSJOED. ISSN 0889-9932.

Anonymous:1997:VPS

[Ano97d] Anonymous. The visual programmer (shepherd and wingo). Microsoft Systems Journal, 12(12): 77–??, December 1997. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1997:WCS

[Ano97e] Anonymous. Windows CE SDK.

MSN 2.0: The inside story. tooltips. Microsoft Systems Journal, 12(4):??, April 1997. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1998:B

[Ano98a] Anonymous. Bugslayer. Microsoft Systems Journal, 13(2): 87–??, February 1998. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1998:C

[Ano98b] Anonymous.  C++.  Mi-

crosoft Systems Journal, 13(2): 101–??, February 1998. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1998:UH

[Ano98c] Anonymous. Under the hood. Microsoft Systems Journal, 13(2): 73–??, February 1998. CODEN MSJOED. ISSN 0889-9932.

Anonymous:1998:VP

[Ano98d] Anonymous. The visual programmer. Microsoft Systems Journal, 13(2):97–??, February 1998. CO-DEN MSJOED. ISSN 0889-9932.

Armstrong:1996:PCC

[Arm96a] Strohm Armstrong. Previewing the Common Controls DLL for Microsoft Internet Explorer 4.0, part II. Microsoft Systems Journal, 11(11):19–??, November 1996. CODEN MSJOED. ISSN 0889-9932.

Armstrong:1996:PNI

[Arm96b] Strohm Armstrong. Previewing the new and improved Common Controls DLL for Microsoft IE 4.0. Microsoft Systems Journal, 11(11):53–??, October 1996. CO-DEN MSJOED. ISSN 0889-9932.

Allard:1995:GWI

[AS95] J. Allard and S. Sinofsky. Getting wired into the Internet: A crash course on FTP, Gopher, Web, and more. Microsoft Systems Journal, 10(5):1–??, May 1995. CODEN MSJOED. ISSN 0889-9932.

Aziz:1994:SEY

[Azi94] Atif Aziz. Simplify and enhance your application’s user interface with dynamic dialog boxes. Microsoft Systems Journal, 9(3): 39–??, March 1994. CODEN MSJOED. ISSN 0889-9932.

Baker:1993:ONP

[Bak93] Steven Baker. An overview of network programming interfaces for Windows and Windows NT. Microsoft Systems Journal, 8(11): 61–??, November 1993. CODEN MSJOED. ISSN 0889-9932.

Box:1998:ECP

[BBES98] Don Box, Keith Brown, Timothy J. Ewald, and Chris Sells. Effective COM programming: Seven tips for building better COMbased applications. Microsoft Systems Journal, 13(10):47–??, October 1998. CODEN MSJOED. ISSN 0889-9932.

Bishop:1998:SMT

[BBM98] F. Avery Bishop, David C. Brown, and David M. Meltzer. Supporting multilanguage text layout and complex scripts with Windows NT 5.0. Microsoft Systems Journal, 13 (11):57–??, November 1998. CO-DEN MSJOED. ISSN 0889-9932.

Bukovec:1998:UMM

[BD98] Mark Bukovec and Dick Dievendorff. Use MSMQ and MTS to simplify the building of transactional applications. Microsoft Systems Journal, 13(7):27–??, July 1998. CODEN MSJOED. ISSN 0889-9932.

Bergman:1990:MCV

[Ber90] Noel J. Bergman. Microsoft C version 6.0 provides an integrated development environment. Microsoft Systems Journal, 5(2): 53–??, March 1990. CODEN MSJOED. ISSN 0889-9932.

Bilson:1992:AFT

[Bil92] Gretchen Bilson. Adapting formal testing techniques for Windows programs. Microsoft Systems Journal, 7(1):77–??, January 1992. CODEN MSJOED. ISSN 0889-9932.

Blaszczak:1996:WIW

[Bla96] Mike Blaszczak. Writing interactive Web apps is a piece of cake with the new ISAPI classes in MFC 4.1. Microsoft Systems Journal, 11(5):41–??, May 1996. CO-DEN MSJOED. ISSN 0889-9932.

Boling:1992:STW

[Bol92] Douglas Boling. Strategies and techniques for writing state-of-the-art TSRs that exploit MS-DOS 5. Microsoft Systems Journal, 7(1):41–??, January 1992. CODEN MSJOED. ISSN 0889-9932.

Boling:1998:MMF

[Bol98] Douglas Boling. Minimizing the memory footprint of your Windows CE-based program. Microsoft Systems Journal, 13(5):37– ??, May 1998. CODEN MSJOED. ISSN 0889-9932.

Box:1995:OQa

[Box95a] Don Box. OLE Q and A. Microsoft Systems Journal, 10(8): 77–??, August 1995. CODEN MSJOED. ISSN 0889-9932.

Box:1995:OQc

[Box95b] Don Box. OLE Q and A. Microsoft Systems Journal, 10(10): 91–??, October 1995. CODEN MSJOED. ISSN 0889-9932.

Box:1995:OQb

[Box95c] Don Box. OLE Q and A. Microsoft Systems Journal, 10(12): 83–??, December 1995. CODEN MSJOED. ISSN 0889-9932.

Box:1995:OQQ

[Box95d] Don Box. OLE Q and Q. Microsoft Systems Journal, 10(6):71– ??, June 1995. CODEN MSJOED. ISSN 0889-9932.

Box:1996:AXC

[Box96a] Don Box. Active X/com Q and A. Microsoft Systems Journal, 11 (11):91–??, November 1996. CO-DEN MSJOED. ISSN 0889-9932.

Box:1996:IDC

[Box96b] Don Box. Introducing distributed COM and the new OLE features in Windows NTM 4.0. Microsoft Systems Journal, 11(5):19– ??, May 1996. CODEN MSJOED. ISSN 0889-9932.

Box:1996:OQAa

[Box96c] Don Box. OLE Q and A. Microsoft Systems Journal, 11(2):

99–??, February 1996. CODEN MSJOED. ISSN 0889-9932.

Box:1996:OQAb

[Box96d] Don Box. OLE Q and A. Microsoft Systems Journal, 11(4): 89–??, April 1996. CODEN MSJOED. ISSN 0889-9932.

Box:1996:OQa

[Box96e] Don Box. OLE Q and A. Microsoft Systems Journal, 11(6):85– ??, June 1996. CODEN MSJOED. ISSN 0889-9932.

Box:1996:OQb

[Box96f] Don Box. OLE Q and A. Microsoft Systems Journal, 11(9):97– ??, September 1996. CODEN MSJOED. ISSN 0889-9932.

Box:1997:ATL

[Box97a] Don Box. The active template library makes building compact COM ObJects a joy. Microsoft Systems Journal, 12(6):41– ??, June 1997. CODEN MSJOED. ISSN 0889-9932.

Box:1997:AXC

[Box97b] Don Box. Active X/COM. Microsoft Systems Journal, 12(11): 73–??, November 1997. CODEN MSJOED. ISSN 0889-9932.

Box:1997:ACa

[Box97c] Don Box. ActiveX/COM. Microsoft Systems Journal, 12(3): 93–??, March 1997. CODEN MSJOED. ISSN 0889-9932.

Box:1997:ACb

[Box97d] Don Box. ActiveX/ com. Microsoft Systems Journal, 12(5):95– ??, May 1997. CODEN MSJOED. ISSN 0889-9932.

Box:1997:ACc

[Box97e] Don Box. ActiveX/COM. Microsoft Systems Journal, 12(9):97– ??, September 1997. CODEN MSJOED. ISSN 0889-9932.

Box:1998:ACa

[Box98a] Don Box. ActiveX/ com. Microsoft Systems Journal, 13(1): 89–??, January 1998. CODEN MSJOED. ISSN 0889-9932.

Box:1998:ACb

[Box98b] Don Box. ActiveX/COM. Microsoft Systems Journal, 13(3): 89–??, March 1998. CODEN MSJOED. ISSN 0889-9932.

Box:1998:ACc

[Box98c] Don Box. ActiveX/ com. Microsoft Systems Journal, 13(5):81– ??, May 1998. CODEN MSJOED. ISSN 0889-9932.

Box:1998:HC

[Box98d] Don Box. House of COM. Microsoft Systems Journal, 13(7):73– ??, July 1998. CODEN MSJOED. ISSN 0889-9932.

Braginski:1998:TIC

[BP98] Leon Braginski and Matt Powell.

Take IIS customization to the next level by writing ISAPI filters and script interpreters. Microsoft Systems Journal, 13(4):51–??, April 1998. CODEN MSJOED. ISSN 0889-9932.

Brockschmidt:1992:PDC

[Bro92] Kraig Brockschmidt. A primer on designing custom controls. Microsoft Systems Journal, 7(2): 87–??, March 1992. CODEN MSJOED. ISSN 0889-9932.

Brockschmidt:1993:IOPb

[Bro93a] Kraig Brockschmidt. Implementing OLE 2.0, part III: Uniform data transfer with data objects. Microsoft Systems Journal, 8(12): 47–??, December 1993. CODEN MSJOED. ISSN 0889-9932.

Brockschmidt:1993:IOPa

[Bro93b] Kraig Brockschmidt. Introducing OLE 2.0, part I: Windows objects and the component object model. Microsoft Systems Journal, 8(8):15–??, August 1993. CO-DEN MSJOED. ISSN 0889-9932.

Brockschmidt:1993:NDW

[Bro93c] Kraig Brockschmidt. Network DDE in Windows for workgroups 3.1 bridges programs between PCs. Microsoft Systems Journal, 8 (1):61–??, January 1993. CODEN MSJOED. ISSN 0889-9932.

Brockschmidt:1993:OPI

[Bro93d] Kraig Brockschmidt. OLE 2.0, part II: Implementing a simple Windows object using either C or C++. Microsoft Systems Journal, 8(9):49–??, September 1993. CO-DEN MSJOED. ISSN 0889-9932.

Brockschmidt:1995:DCI

[Bro95]  Kraig Brockschmidt.  Design

considerations for implementing a simple OLE automation controller. Microsoft Systems Journal, 10(5):57–??, May 1995. CO-DEN MSJOED. ISSN 0889-9932.

Brockschmidt:1996:HCS

[Bro96a] Kraig Brockschmidt. How COM solves the problems of component software design, part II. Microsoft Systems Journal, 11(6):19– ??, June 1996. CODEN MSJOED. ISSN 0889-9932.

Brockschmidt:1996:HOC

[Bro96b] Kraig Brockschmidt. How OLE and COM solve the problems of component software design. Microsoft Systems Journal, 11(5):63– ??, May 1996. CODEN MSJOED. ISSN 0889-9932.

Campbell:1996:EWE

[Cam96] David Campbell. Extending the Windows Explorer with name space extensions. Microsoft Systems Journal, 11(7):41–??, July 1996. CODEN MSJOED. ISSN 0889-9932.

Campbell:1997:HEM

[Cam97] David Campbell. How to exploit multiple monitor support in Memphis and Windows NT 5.0. Microsoft Systems Journal, 12(6):17– ??, June 1997. CODEN MSJOED. ISSN 0889-9932.

Capucciati:1993:PYB

[Cap93] Maria R. Capucciati. Putting your best face forward: Designing an effective user interface. Microsoft Systems Journal, 8(2): 55–??, February 1993. CODEN MSJOED. ISSN 0889-9932.

Cluts:1997:PAC

[CE97] Nancy Winnick Cluts and Michael Edwards. A preview of Active Channel and the Active Desktop for Internet Explorer 4.0. Microsoft Systems Journal, 12(10): 59–??, October 1997. CODEN MSJOED. ISSN 0889-9932.

Chappell:1998:HMT

[Cha98a] David Chappell. How Microsoft Transaction Server changes the COM programming model. Microsoft Systems Journal, 13(1): 19–??, January 1998. CODEN MSJOED. ISSN 0889-9932.

Chappell:1998:MMQ

[Cha98b] David Chappell. Microsoft Message Queue is a fast, efficient choice for your distributed application. Microsoft Systems Journal, 13(7):17–??, July 1998. CO-DEN MSJOED. ISSN 0889-9932.

Chiverton:1990:MYM

[Chi90] Bob Chiverton. Measure your machine’s activity and learn how to use OS/2 threads with CPUME-TER. Microsoft Systems Journal, 5(6):49–??, November 1990. CO-DEN MSJOED. ISSN 0889-9932.

Chiverton:1991:PHW

[Chi91] Bob Chiverton. Postman, a handy Windows debugging tool, lets you post messages to any window. Microsoft Systems Journal, 6(3):21– ??, May 1991. CODEN MSJOED. ISSN 0889-9932.

Chiverton:1992:SSL

[Chi92]  Bob Chiverton.  Shed some

light on your Windows application’s default data segment with HeapPeep. Microsoft Systems Journal, 7(1):17–??, January 1992. CODEN MSJOED. ISSN 0889-9932.

Chiverton:1993:VCT

[Chi93] Bob Chiverton. Visual C++ enters the ring swinging and scores a technical knockout. Microsoft Systems Journal, 8(6):15–??, June 1993. CODEN MSJOED. ISSN 0889-9932.

Chiverton:1994:VCT

[Chi94] Bob Chiverton. Visual C++ 2.0 features better editors, faster builds, and C++ exceptions and templates. Microsoft Systems Journal, 9(7):37–??, July 1994. CODEN MSJOED. ISSN 0889-9932.

Chiverton:1995:MDO

[Chi95] Bob Chiverton. MFC 4.0 delivers OCX containers, improved DLL models, and new classes. Microsoft Systems Journal, 10(12): 19–??, December 1995. CODEN MSJOED. ISSN 0889-9932.

Clark:1998:EDP

[Cla98a] Jason Clark. Exploring DirectX 5.0, part II: DirectSound gives your apps advanced 3D sound. Microsoft Systems Journal, 13 (6):21–??, June 1998. CODEN MSJOED. ISSN 0889-9932.

Clark:1998:MFF

[Cla98b] Jason Clark. May the force feedback be with you: Grappling with DirectX and DirectInput. Microsoft Systems Journal, 13(2): 19–??, February 1998. CODEN MSJOED. ISSN 0889-9932.

Comeau:1990:FRC

[Com90] Greg Comeau. Fundamental recommendations on C programming style. Microsoft Systems Journal, 5(3):67–??, May 1990. CODEN MSJOED. ISSN 0889-9932.

Cook:1996:WSH

[Coo96] David Cook. Write a simple HTTP based server using MFC and Windows sockets. Microsoft Systems Journal, 11(2): 43–??, February 1996. CODEN MSJOED. ISSN 0889-9932.

Cunderlik:1998:TAM

[Cun98] Maros Cunderlik. Take advantage of MTS in your distributed system with custom resource dispensers. Microsoft Systems Journal, 13(9): 29–??, September 1998. CODEN MSJOED. ISSN 0889-9932.

Custer:1992:GTW

[Cus92] Helen Custer. A grand tour of Windows NT: Portable 32bit multiprocessing comes to windows. Microsoft Systems Journal, 7(4):17–??, July 1992. CODEN MSJOED. ISSN 0889-9932.

Custer:1995:DID

[Cus95] Helen Custer. Design and implementation details of the Windows NT virtual block cache manager. Microsoft Systems Journal, 10(7):67–??, July 1995. CODEN MSJOED. ISSN 0889-9932.

Daniell:1998:IWV

[DDBM98] Todd Daniell, Brian Daigle, Doug Bahr, and Dave Mims. Implementing a Web view namespace extension using active directory services. Microsoft Systems Journal, 13(8):37–??, August 1998. CODEN MSJOED. ISSN 0889-9932.

DeLascia:1998:C

[DeL98] Paul DeLascia. C++. Microsoft Systems Journal, 13(1): 103–??, January 1998. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1993:CCQ

[DiL93a] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 8 (5):77–??, May 1993. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1993:FMT

[DiL93b] Paul DiLascia. Fun with MFC: 33 tips to help you get the most out of C++ class libraries for Windows. Microsoft Systems Journal, 8(11): 17–??, November 1993. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1994:CCQa

[DiL94a] Paul DiLascia. C/C++ Q and A.

Microsoft Systems Journal, 9(2): 85–??, February 1994. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1994:CCQb

[DiL94b] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 9 (5):71–??, May 1994. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1994:OMA

[DiL94c] Paul DiLascia. OLE made almost easy: Creating containers and servers using MFC 2.5. Microsoft Systems Journal, 9(4): 13–??, April 1994. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1995:CCQd

[DiL95a] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 10 (5):93–??, May 1995. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1995:CCQa

[DiL95b] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 10 (6):87–??, June 1995. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1995:CCQg

[DiL95c] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 10(7):99–??, July 1995. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1995:CCQe

[DiL95d] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 10 (8):99–??, August 1995. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1995:CCQc

[DiL95e] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 10(9): 89–??, September 1995. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1995:CCQh

[DiL95f] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 10 (10):99–??, October 1995. CO-DEN MSJOED. ISSN 0889-9932.

DiLascia:1995:CCQb

[DiL95g] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 10 (11):89–??, November 1995. CO-DEN MSJOED. ISSN 0889-9932.

DiLascia:1995:CCQf

[DiL95h] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 10 (12):99–??, December 1995. CO-DEN MSJOED. ISSN 0889-9932.

DiLascia:1995:MTM

[DiL95i] Paul DiLascia. Meandering through the maze of MFC message and command routing. Microsoft Systems Journal, 10(7):17– ??, July 1995. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1996:CQa

[DiL96a] Paul DiLascia. C++ Q and A. Microsoft Systems Journal, 11(7):91– ??, July 1996. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1996:CQb

[DiL96b] Paul DiLascia. C++ Q and A.

Microsoft Systems Journal, 11(8): 87–??, August 1996. CODEN MSJOED. ISSN 0889-9932.

Dilascia:1996:C

[Dil96c] Paul Dilascia. C++ Q and A.

Microsoft Systems Journal, 11(9): 103–??, September 1996. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1996:CQc

[DiL96d] Paul DiLascia. C++ Q and A. Microsoft Systems Journal, 11(12): 89–??, December 1996. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1996:CCQa

[DiL96e] Paul DiLascia. C/C++ Q and A.

Microsoft Systems Journal, 11(1): 103–??, January 1996. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1996:CCQb

[DiL96f] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 11(2): 107–??, February 1996. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1996:CCQc

[DiL96g] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 11 (4):107–??, April 1996. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1996:CCQd

[DiL96h] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 11 (5):99–??, May 1996. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1996:CCQe

[DiL96i] Paul DiLascia. C/C++ Q and A. Microsoft Systems Journal, 11 (6):99–??, June 1996. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1996:CC

[DiL96j] Paul DiLascia. C/C++ Q+A. Microsoft Systems Journal, 11(3): 103–??, March 1996. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1997:Cb

[DiL97a] Paul DiLascia. C++. Microsoft Systems Journal, 12(12): 103–??, December 1997. CODEN MSJOED. ISSN 0889-9932.

Dilascia:1997:CQa

[Dil97b] Paul Dilascia. C++ Q and A. Microsoft Systems Journal, 12(2): 101–??, February 1997. CODEN MSJOED. ISSN 0889-9932.

Dilascia:1997:CQb

[Dil97c] Paul Dilascia. C++ Q and A.

Microsoft Systems Journal, 12(3): 109–??, March 1997. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1997:CQc

[DiL97d] Paul DiLascia. C++ Q and A. Microsoft Systems Journal, 12 (4):93–??, April 1997. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1997:CQd

[DiL97e] Paul DiLascia. C++ Q and A. Microsoft Systems Journal, 12 (5):113–??, May 1997. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1997:CQe

[DiL97f] Paul DiLascia. C++ Q and A.

Microsoft Systems Journal, 12(8): 107–??, August 1997. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1997:CQf

[DiL97g] Paul DiLascia. C++ Q and A.

Microsoft Systems Journal, 12(9): 109–??, September 1997. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1997:Ca

[DiL97h] Paul DiLascia. C++A and A.

Microsoft Systems Journal, 12(6): 109–??, June 1997. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1997:MFMb

[DiL97i] Paul DiLascia. More fun with MFC: DIBs, palettes, subclassing, and a gamut of goodies, part II. Microsoft Systems Journal, 12 (3):47–??, March 1997. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1997:MFMc

[DiL97j] Paul DiLascia. More fun with MFC: DIBs, palettes, subclassing, and a gamut of goodies, part III. Microsoft Systems Journal, 12(6):69–??, June 1997. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1997:MFMa

[DiL97k] Paul DiLascia. More fun with MFC: DIBs, palettes, subclassing, and a gamut of reusable goodies. Microsoft Systems Journal, 12(1): 17–??, January 1997. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1998:Ca

[DiL98a] Paul DiLascia. C++. Microsoft Systems Journal, 13(3): 99–??, March 1998. CODEN MSJOED. ISSN 0889-9932.

Dilascia:1998:C

[Dil98b] Paul Dilascia.  C++.  Mi-

crosoft Systems Journal, 13(4): 103–??, April 1998. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1998:Cb

[DiL98c] Paul DiLascia. C++. Microsoft Systems Journal, 13(5): 105–??, May 1998. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1998:Cc

[DiL98d] Paul DiLascia. C++. Microsoft Systems Journal, 13(6):11– ??, June 1998. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1998:Cd

[DiL98e] Paul DiLascia. C++. Microsoft Systems Journal, 13(7):83– ??, July 1998. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1998:GYA

[DiL98f] Paul DiLascia. Give your applications the hot new interface look with cool menu buttons. Microsoft Systems Journal, 13(1): 31–??, January 1998. CODEN MSJOED. ISSN 0889-9932.

DiLascia:1998:VCB

[DiL98g] Paul DiLascia. Visual C++6.0 brings home a full bag of tricks and treats for MFC developers. Microsoft Systems Journal, 13 (10):29–??, October 1998. CO-DEN MSJOED. ISSN 0889-9932.

Dixon:1991:CCA

[Dix91a] Brendan W. Dixon. Creating a client-server application with LAN Manager 2.0 and named pipes. Microsoft Systems Journal, 6(2):93–??, March 1991. CODEN MSJOED. ISSN 0889-9932.

Dixon:1991:CNS

[Dix91b] Brendan W. Dixon. Creating a network service using the clientserver model and LAN Manager 2.0. Microsoft Systems Journal, 6 (1):45–??, January 1991. CODEN MSJOED. ISSN 0889-9932.

Duncan:1990:OVE

[Dun90a] Ray Duncan. OS/2 version 2.0:

Exploiting the 32-bit architecture of 80386- and 80486-based systems. Microsoft Systems Journal, 5(3):01–??, May 1990. CODEN MSJOED. ISSN 0889-9932.

Duncan:1990:SPA

[Dun90b] Ray Duncan. Supporting protected-mode applications in a DOS-based environment. Microsoft Systems Journal, 5(4):92– ??, July 1990. CODEN MSJOED. ISSN 0889-9932.

Edson:1992:DTT

[Eds92a] Dave Edson. Dave’s top ten list of tricks, hints, and techniques for programming in Windows. Microsoft Systems Journal, 7(6):31–??, October 1992. CO-DEN MSJOED. ISSN 0889-9932.

Edson:1992:HTI

[Eds92b] Dave Edson. How to track, isolate, and exterminate bugs in your Windows-based applications. Microsoft Systems Journal, 7(3):69– ??, May 1992. CODEN MSJOED. ISSN 0889-9932.

Edson:1993:SPB

[Eds93] Dave Edson. Slay the porting beasties: Dave’s top ten tips for migrating to Windows NT. Microsoft Systems Journal, 8(9):37– ??, September 1993. CODEN MSJOED. ISSN 0889-9932.

Edson:1994:CCC

[Eds94a] Dave Edson. Cleverly coding with Chicago’s interface gadgets, part I: Image lists and TreeViews. Microsoft Systems Journal, 9(7):13– ??, July 1994. CODEN MSJOED. ISSN 0889-9932.

Edson:1994:GBW

[Eds94b] Dave Edson. Going beyond Word-Basic: Extending word 6.0 for Windows with WLLs. Microsoft Systems Journal, 9(5):41–??, May 1994. CODEN MSJOED. ISSN 0889-9932.

Edson:1994:STP

[Eds94c] Dave Edson. Seventeen techniques for preparing your 16-bit applications for Chicago. Microsoft Systems Journal, 9(2):15–??, February 1994. CODEN MSJOED. ISSN 0889-9932.

Edson:1995:GSW

[Eds95a] Dave Edson. The game SDK for Windows 95 gives you direct hardware control for speedy animation. Microsoft Systems Journal, 10(11):15–??, November 1995. CODEN MSJOED. ISSN 0889-9932.

Edson:1995:GBG

[Eds95b] Dave Edson. A grab bag of gotchas and goodies for programming in Windows 95. Microsoft Systems Journal, 10(5):19– ??, May 1995. CODEN MSJOED. ISSN 0889-9932.

Edson:1996:GWN

[Eds96] Dave Edson. Get world-class noise and total joy from your games with direct and directinput. Microsoft Systems Journal, 11(2): 71–??, February 1996. CODEN MSJOED. ISSN 0889-9932.

Eddon:1995:PYW

[EE95] Guy Eddon and Henry Eddon.

Port your Windows-based app to the Mac with the Visual C++ cross-development edition. Microsoft Systems Journal, 10(6):53– ??, June 1995. CODEN MSJOED. ISSN 0889-9932.

Eddon:1996:ECS

[EE96] Guy Eddon and Henry Eddon. Exploring the clien/server capabilities of Visual Basic 4.0 enterprise edition. Microsoft Systems Journal, 11(3):55–??, March 1996. CODEN MSJOED. ISSN 0889-9932.

Eddon:1997:VBRa

[EE97a] Guy Eddon and Henry Eddon. Visual Basic 5.0 relieves the pain and discomfort of ActiveX control creation. Microsoft Systems Journal, 12(2):57–??, February 1997. CO-DEN MSJOED. ISSN 0889-9932.

Eddon:1997:VBRb

[EE97b] Guy Eddon and Henry Eddon. Visual Basic 5.0 relieves the pain and discomfort of ActiveX control creation, part II. Microsoft Systems Journal, 12(3):37–??, March 1997. CODEN MSJOED. ISSN 0889-9932.

Eddon:1998:UDW

[EE98] Guy Eddon and Henry Eddon. Understanding the DCOM wire protocol by analyzing network data packets. Microsoft Systems Journal, 13(3):45–??, March 1998. CODEN MSJOED. ISSN 0889-9932.

Eisner:1990:EOW

[Eis90] Bill Eisner. Extending off-the-shelf Windows applications with macros and DLLs. Microsoft Systems Journal, 5(6):89–??, November 1990. CODEN MSJOED. ISSN 0889-9932.

Euler:1991:CRW

[EMR91] Laura Euler, Eric Maffei, and Adam Rauch. Create real Windows applications in a graphical environment using Microsoft Visual Basic. Microsoft Systems Journal, 6(4):57–??, July 1991. CODEN MSJOED. ISSN 0889-9932.

Faison:1994:PBS

[Fai94a] Ted Faison. A primer on building server-style DLLs for Windows with MFC 2.5 and OWL 2.0. Microsoft Systems Journal, 9(6):43– ??, June 1994. CODEN MSJOED. ISSN 0889-9932.

Faison:1994:POC

[Fai94b] Ted Faison. Putting the OWL 2.0 class library for Windows through its paces. Microsoft Systems Journal, 9(2):45–??, February 1994. CODEN MSJOED. ISSN 0889-9932.

Faison:1998:NHH

[Fai98] Ted Faison. The new HTML help system extends online help from the desktop to the web. Microsoft Systems Journal, 13(1): 55–??, January 1998. CODEN MSJOED. ISSN 0889-9932.

Finnegan:1993:WTQ

[Fin93] Fran Finnegan. Windows Q and A. Microsoft Systems Journal, 8 (5):67–??, May 1993. CODEN MSJOED. ISSN 0889-9932.

Finnegan:1994:TDW

[Fin94]  James Finnegan.  Test drive

Win32 from 16-bit code using the Windows NT WOW layer and generic thunk. Microsoft Systems Journal, 9(6):13–??, June 1994. CODEN MSJOED. ISSN 0889-9932.

Finnegan:1995:BWN

[Fin95a] J. Finnegan. Building Windows NT-Based client/server applications using remote procedure calls. Microsoft Systems Journal, 10 (5):43–??, May 1995. CODEN MSJOED. ISSN 0889-9932.

Finnegan:1995:DCS

[Fin95b] James Finnegan. Developing a client/server multimedia app using RPCs and video for Windows 1.1. Microsoft Systems Journal, 10 (9):53–??, September 1995. CO-DEN MSJOED. ISSN 0889-9932.

Finnegan:1995:HMW

[Fin95c] James Finnegan. Hook and monitor any 16-bit Windows function with our ProcHook DLL. Microsoft Systems Journal, 10(1): 45–??, January 1995. CODEN MSJOED. ISSN 0889-9932.

Finnegan:1998:POP

[Fin98a] James Finnegan. Pop open a privileged set of APls with Windows NT kernel mode drivers. Microsoft Systems Journal, 13(3): 19–??, March 1998. CODEN MSJOED. ISSN 0889-9932.

Finocchio:1998:CSEa

[Fin98b] Mark Finocchio. Control Spy exposes the clandestine life of Windows Common Controls, part I. Microsoft Systems Journal, 13 (7):43–??, July 1998. CODEN MSJOED. ISSN 0889-9932.

Finocchio:1998:CSEb

[Fin98c] Mark Finocchio. Control Spy exposes the clandestine life of Windows Common Controls. part II. Microsoft Systems Journal, 13(9): 51–??, September 1998. CODEN MSJOED. ISSN 0889-9932.

Freytag:1990:UON

[FL90] Asmus Freytag and Michael Leu.

Using the OS/2 national language support services to write international programs. Microsoft Systems Journal, 5(2):01–??, March 1990. CODEN MSJOED. ISSN 0889-9932.

Fosner:1996:OPC

[Fos96] Ron Fosner. OpenGL without the pain: Creating a reusable 3D view class for MFC. Microsoft Systems Journal, 11(11):41–??, November 1996. CODEN MSJOED. ISSN 0889-9932.

Fosner:1998:GFS

[Fos98] Ron Fosner. Get fast and simple 3D rendering with draw primitive and DirectX 5.0. Microsoft Systems Journal, 13(2):37–??, February 1998. CODEN MSJOED. ISSN 0889-9932.

Fishman:1997:WCS

[FR97] Neil Fishman and Jeffrey Richter.

The Windows CE SDK: The tools you need to program the handheld PC. Microsoft Systems Journal, 12(4):17–??, April 1997. CODEN MSJOED. ISSN 0889-9932.

Fogelin:1990:PWT

[FWB90] Eric Fogelin, David Wood, and Noel Bergman. Port Windows applications to OS/2 (almost) painlessly with the software migration kit. Microsoft Systems Journal, 5(6):21–??, November 1990. CODEN MSJOED. ISSN 0889-9932.

Geary:1990:IMW

[Gea90] Michael Geary. An introduction to Microsoft Windows version 3.0: A developer’s viewpoint. Microsoft Systems Journal, 5(4):01–??, July 1990. CODEN MSJOED. ISSN 0889-9932.

Gellock:1993:PYA

[Gel93] Scot Gellock. Port your 16bit applications to Windows NT without ripping your hair out. Microsoft Systems Journal, 8(8): 47–??, August 1993. CODEN MSJOED. ISSN 0889-9932.

Goodman:1992:CEM

[Goo92] Kevin J. Goodman. A comprehensive examination of the Microsoft C version 6.0 startup code. Microsoft Systems Journal, 7(1): 63–??, January 1992. CODEN MSJOED. ISSN 0889-9932.

Goodman:1995:CMC

[Goo95] Kevin Goodman. Clearer, more comprehensive error processing with Win32 structured exception handling. Microsoft Systems Journal, 10(1):29–??, January 1995. CODEN MSJOED. ISSN 0889-9932.

Greenberg:1990:APP

[Gre90] Ross M. Greenberg. Adapting the parallel port for bidirectional communication. Microsoft Systems Journal, 5(5):107–??, September 1990. CODEN MSJOED. ISSN 0889-9932.

Hall:1991:AYP

[Hal91a] William Hall. Adapt your program for worldwide use with Windows internationalization support. Microsoft Systems Journal, 6(6): 29–??, November 1991. CODEN MSJOED. ISSN 0889-9932.

Hall:1991:AEP

[Hal91b] William Hall. Adapting extended processes to the cooperative multitasking of Microsoft Windows. Microsoft Systems Journal, 6(1): 21–??, January 1991. CODEN MSJOED. ISSN 0889-9932.

Hall:1994:IWNa

[Hal94a] William S. Hall. Internationalization in Windows NT, part I: Programming with Unicode. Microsoft Systems Journal, 9(6):57– ??, June 1994. CODEN MSJOED. ISSN 0889-9932.

Hall:1994:IWNb

[Hal94b] William S. Hall. Internationalization in Windows NT, part II: Locales, languages, and resources. Microsoft Systems Journal, 9(7):55–??, July 1994. CO-DEN MSJOED. ISSN 0889-9932.

Heller:1995:TTD

[Hel95] Martin Heller. Tips and tricks on developing killer server applications for Windows NT. Microsoft Systems Journal, 10(8):41–??, August 1995. CODEN MSJOED. ISSN 0889-9932.

Hildebrand:1990:TDT

[Hil90] Daniel Hildebrand. TRACER: A debugging tool for OS/2 Presentation Manager development. Microsoft Systems Journal, 5(2): 63–??, March 1990. CODEN MSJOED. ISSN 0889-9932.

Hludzinski:1998:UID

[Hlu98] Bill Hludzinski. Understanding interface definition language: A developer’s survival guide. Microsoft Systems Journal, 13(8):51–??, August 1998. CODEN MSJOED. ISSN 0889-9932.

Holub:1993:VCT

[Hol93] Allen Holub. Visual C++: Its compiler, language implementation, and code quality. Microsoft Systems Journal, 8(6):65–??, June 1993. CODEN MSJOED. ISSN 0889-9932.

Holub:1995:RMS

[Hol95] Allen Holub. Rewriting the MFC scribble program using an objectoriented design approach. Microsoft Systems Journal, 10(8): 17–??, August 1995. CODEN MSJOED. ISSN 0889-9932.

Holub:1996:HYDa

[Hol96a] Allen Holub. Hide your data and make objects responsible for their own user interfaces. Microsoft Systems Journal, 11(8):53–??, August 1996. CODEN MSJOED. ISSN 0889-9932.

Holub:1996:HYDb

[Hol96b] Allen Holub. Hide your data and make objects responsible for their own user interfaces, part II. Microsoft Systems Journal, 11(12): 35–??, December 1996. CODEN MSJOED. ISSN 0889-9932.

Holub:1996:RYO

[Hol96c] Allen Holub. Roll your own persistence implementations to go beyond the MFC frontier. Microsoft Systems Journal, 11(6):31– ??, June 1996. CODEN MSJOED. ISSN 0889-9932.

Holub:1997:HYD

[Hol97] Allen Holub. Hide your data and make objects responsible for their own user interfaces, part III. Microsoft Systems Journal, 12(1): 67–??, January 1997. CODEN MSJOED. ISSN 0889-9932.

Huckins:1996:OEC

[Huc96] Jeff Huckins. Our exclusive class library speeds building Windows NT kernel-mode device drivers. Microsoft Systems Journal, 11(9): 35–??, September 1996. CODEN MSJOED. ISSN 0889-9932.

Jack:1997:RTT

[Jac97] Roger Jack. Riptoe through the Tooltips with our allencompassing ToolTip programmers’ guide. Microsoft Systems Journal, 12(4):49–??, April 1997. CODEN MSJOED. ISSN 0889-9932.

Jamsa:1990:CIY

[Jam90a] Kris Jamsa. DOS commands inside your code: Process control and signal handling. Microsoft Systems Journal, 5(1):85–??, January 1990. CODEN MSJOED. ISSN 0889-9932.

Jamsa:1990:EOT

[Jam90b] Kris Jamsa. Examining objectoriented techniques using Quick-Pascal. Microsoft Systems Journal, 5(3):71–??, May 1990. CO-DEN MSJOED. ISSN 0889-9932.

Jongerius:1991:ATW

[Jon91] Jerry Jongerius. Accurately timing Windows events without timer reprogramming. Microsoft Systems Journal, 6(4):75–??, July 1991. CODEN MSJOED. ISSN 0889-9932.

Jones:1998:WCN

[Jon98] Anthony Jones. Windows CE 2.0 networking offers exciting mobile computing possibilities. Microsoft Systems Journal, 13(5):53– ??, May 1998. CODEN MSJOED. ISSN 0889-9932.

Kelly:1998:GCA

[Kel98] Mike Kelly. Gain control of application setup and maintenance with the new Windows Installer. Microsoft Systems Journal, 13(9): 15–??, September 1998. CODEN MSJOED. ISSN 0889-9932.

Kerber:1990:SCW

[Ker90]  Scott Kerber.  Simplifying

complex Windows development through the use of a clientserver architecture. Microsoft Systems Journal, 5(1):21–??, January 1990. CODEN MSJOED. ISSN 0889-9932.

Keyser:1992:EFW

[Key92] Greg Keyser. Extend the functionality of the Windows control panel with custom DLLs. Microsoft Systems Journal, 7(7):45–??, November 1992. CODEN MSJOED. ISSN 0889-9932.

Kim:1998:WDC

[Kim98] Frank Kim. Why do certain Win32 technologies misbehave in Windows NT services? Microsoft Systems Journal, 13(3): 67–??, March 1998. CODEN MSJOED. ISSN 0889-9932.

King:1994:MMW

[Kin94] Adrian King. Memory management, the Win32 subsystem, and internal synchronization in Chicago. Microsoft Systems Journal, 9(5):57–??, May 1994. CO-DEN MSJOED. ISSN 0889-9932.

King:1995:WST

[Kin95] Adrian King. Windows, the next generation: An advance look at the architecture of Chicago. Microsoft Systems Journal, 10(1): 15–??, January 1995. CODEN MSJOED. ISSN 0889-9932.

Kirtland:1996:SWS

[Kir96] Mary Kirtland. Safe Web surfing with the Internet component download service. Microsoft Systems Journal, 11(7):65–??, July 1996. CODEN MSJOED. ISSN 0889-9932.

Kirtland:1997:OOS

[Kir97a] Marty Kirtland. Object oriented software development made simple with COM+ runtime services. Microsoft Systems Journal, 12 (11):49–??, November 1997. CO-DEN MSJOED. ISSN 0889-9932.

Kirtland:1997:CPM

[Kir97b] Mary Kirtland. The COM+ programming model makes it easy to write components in any language. Microsoft Systems Journal, 12(12):19–??, December 1997. CODEN MSJOED. ISSN 0889-9932.

Kirtland:1997:IVS

[Kir97c] Mary Kirtland. Introducing Visual Studio 97: A well stocked toolbox for building distributed apps. Microsoft Systems Journal, 12(5):19–??, May 1997. CODEN MSJOED. ISSN 0889-9932.

Klemond:1992:IOL

[Kle92a] Paul Klemond. Investigating object linking and embedding, part II: Adding server support. Microsoft Systems Journal, 7(3):87– ??, May 1992. CODEN MSJOED. ISSN 0889-9932.

Klemond:1992:TBH

[Kle92b] Paul Klemond. Taking the bull by the horns: Investigating object linking and embedding, part I. Microsoft Systems Journal, 7 (2):19–??, March 1992. CODEN MSJOED. ISSN 0889-9932.

Lang:1995:BCS

[Lan95] E. Lang. Building component software with Visual C++ and the OLE custom control developer’s kit. Microsoft Systems Journal, 10(5):25–??, May 1995. CODEN MSJOED. ISSN 0889-9932.

Massoni:1998:ILA

[Mas98] Joe Massoni. An inside look at developing applications using the new features of visual C++6.0. Microsoft Systems Journal, 13 (10):17–??, October 1998. CO-DEN MSJOED. ISSN 0889-9932.

McKeown:1995:HGY

[McK95] Michael McKeown. How to get your Windows 95 OLE controls refined, robust, and ready to ship. Microsoft Systems Journal, 10(10):67–??, October 1995. CO-DEN MSJOED. ISSN 0889-9932.

McKeown:1996:BOC

[McK96] Michael McKeown. Build OLE controls for the Internet that are fast, smart, and interactive. Microsoft Systems Journal, 11(4): 19–??, April 1996. CODEN MSJOED. ISSN 0889-9932.

Mefford:1990:IEP

[Mef90] Michael J. Mefford. An indepth exploration of the PC keyboard and its interrupt service routines. Microsoft Systems Journal, 5(3):22–??, May 1990. CO-DEN MSJOED. ISSN 0889-9932.

Meyers:1993:BDF

[Mey93a] Brian Meyers. Built-in diagnostic facilities in the Microsoft Foundation Classes simplify debugging. Microsoft Systems Journal, 8(2): 41–??, February 1993. CODEN MSJOED. ISSN 0889-9932.

Meyers:1993:MDF

[Mey93b] Brian Meyers. MFC diagnostic facilities part II: More-sophisticated debugging techniques. Microsoft Systems Journal, 8(3): 55–??, March 1993. CODEN MSJOED. ISSN 0889-9932.

Meyers:1995:HNT

[Mey95] Scott Meyers. How to navigate the treacherous waters of C++ exception handling. Microsoft Systems Journal, 10(11):61–??, November 1995. CODEN MSJOED. ISSN 0889-9932.

Mirho:1991:CFB

[Mir91] C. Albert Mirho. Converting foreign bitmaps to Windows deviceindependent bitmaps. Microsoft Systems Journal, 6(4):85–??, July 1991. CODEN MSJOED. ISSN 0889-9932.

Mirho:1993:IRP

[Mir93]  Charles Mirho.  It’s a raid!

putting four debuggers for Windows through their paces. Microsoft Systems Journal, 8(4): 35–??, April 1993. CODEN MSJOED. ISSN 0889-9932.

Mirho:1996:LAV

[Mir96] Charles Mirho. To learn about the Voice Modem Extensions for Windows(R) 95, press 1 now! Microsoft Systems Journal, 11(8): 39–??, August 1996. CODEN MSJOED. ISSN 0889-9932.

Mitchell:1996:LYV

[Mit96] David C. Mitchell. Leveraging your Visual C++ experience on the Internet with thin client technology. Microsoft Systems Journal, 11(12):47–??, December 1996. CODEN MSJOED. ISSN 0889-9932.

Mirho:1993:RTS

[MR93] Charles Mirho and Andrew Raffman. Reach out and touch someone’s PC: The Windows telephony API. Microsoft Systems Journal, 8(12):15–??, December 1993. CO-DEN MSJOED. ISSN 0889-9932.

Maxey:1991:CSC

[MS91] David Maxey and Andrew Schulman. Call Standard C I/O functions from your Windows code using the WINIO library. Microsoft Systems Journal, 6(4):33–??, July 1991. CODEN MSJOED. ISSN 0889-9932.

Mee:1996:DDU

[MTM96] Michael Mee, Peter Tucker, and John McCullough. The DAO of databases: Using data access objects and the jet engine in C++. Microsoft Systems Journal, 11(1): 35–??, January 1996. CODEN MSJOED. ISSN 0889-9932.

Nicolaisen:1996:VCD

[Nic96] Nancy Nicolaisen. Visual C++ 4.2 dramatically reduces the learning curve for writing Internet apps. Microsoft Systems Journal, 11(9): 69–??, September 1996. CODEN MSJOED. ISSN 0889-9932.

Norton:1992:DAT

[Nor92] Daniel Norton. Developing automated testing procedures with Microsoft test and TestBasic scripts. Microsoft Systems Journal, 7(7): 63–??, November 1992. CODEN MSJOED. ISSN 0889-9932.

Notestein:1996:BMC

[Not96] Jenny Notestein. Building multitiered client/ server apps with Visual Basic 4.0 Enterprise Edition. Microsoft Systems Journal, 11(11):57–??, November 1996.

CODEN MSJOED. ISSN 0889-9932.

Notestein:1997:BMC

[Not97] Jenny Notestein. Building multitiered client/ server apps with Visual Basic 4.0 Enterprise Edition, part II. Microsoft Systems Journal, 12(2):37–??, February 1997. CODEN MSJOED. ISSN 0889-9932.

Opstad:1990:IOC

[OH90] Lars Opstad and Arthur Hanson. Interfacing OS/2 compiled BASIC to Presentation Manager. Microsoft Systems Journal, 5(1): 77–??, January 1990. CODEN MSJOED. ISSN 0889-9932.

Oney:1993:MCY

[One93] Walter Oney. Mix 16-bit and 32-bit code in your applications with the Win32s universal thunk. Microsoft Systems Journal, 8(11): 39–??, November 1993. CODEN MSJOED. ISSN 0889-9932.

Oney:1994:RBY

[One94] Walter Oney. Removing bottlenecks from your program with Windows NT performance-tuning tools. Microsoft Systems Journal, 9(4):35–??, April 1994. CODEN MSJOED. ISSN 0889-9932.

Oney:1995:UFP

[One95a] W. Oney. Unconstrained filenames on the PC! introducing Chicago’s protected mode FAT file system. Microsoft Systems Journal, 10(4):7–??, April 1995. CO-DEN MSJOED. ISSN 0889-9932.

Oney:1995:BDD

[One95b] Walter Oney. Building device drivers that conform to Windows 95. Microsoft Systems Journal, 10 (12):35–??, December 1995. CO-DEN MSJOED. ISSN 0889-9932.

Oney:1995:EYA

[One95c] Walter Oney. Extend your application with dynamically loaded VxDs under Windows 95. Microsoft Systems Journal, 10(5): 133–??, May 1995. CODEN MSJOED. ISSN 0889-9932.

Oney:1996:ACY

[One96] Walter Oney. Automatically configure your devices by exploiting plug and play VxD services. Microsoft Systems Journal, 11(1): 57–??, January 1996. CODEN MSJOED. ISSN 0889-9932.

Oney:1997:INW

[One97a] Walter Oney. Implementing the new Win32 driver model for Windows 98 and Windows NT 5.0. Microsoft Systems Journal, 12(12): 43–??, December 1997. CODEN MSJOED. ISSN 0889-9932.

Oney:1997:PNA

[One97b] Walter Oney. Pay no attention to the man behind the curtain! write your own C++ appwizards. Microsoft Systems Journal, 12(3): 19–??, March 1997. CODEN MSJOED. ISSN 0889-9932.

Oney:1997:SNW

[One97c] Walter Oney. Surveying the new Win32 driver model for Windows 98 and Windows NT 5.0. Microsoft Systems Journal, 12(11):

35–??, November 1997. CODEN MSJOED. ISSN 0889-9932.

Patch:1992:INP

[Pat92] Ray Patch. An introduction to network programming using the NetBIOS interface. Microsoft Systems Journal, 7(2):61–??, March 1992. CODEN MSJOED. ISSN 0889-9932.

Pattison:1998:BMC

[Pat98] Ted Pattison. Build MTS components with visual basic for deployment in your ASP based apps. Microsoft Systems Journal, 13 (4):19–??, April 1998. CODEN MSJOED. ISSN 0889-9932.

Petzold:1990:CPP

[Pet90a] Charles Petzold. Checkers for PM part IV: Adding the game-playing strategy. Microsoft Systems Journal, 5(5):69–??, September 1990. CODEN MSJOED. ISSN 0889-9932.

Petzold:1990:CPM

[Pet90b] Charles Petzold. Checkers for Presentation Manager part III: Moving the pieces. Microsoft Systems Journal, 5(2):75–??, March 1990. CODEN MSJOED. ISSN 0889-9932.

Petzold:1990:CPI

[Pet90c] Charles Petzold. Checkers part II: Drawing the checkerboard and pieces using GPI. Microsoft Systems Journal, 5(1):63–??, January 1990. CODEN MSJOED. ISSN 0889-9932.

Petzold:1990:NMD

[Pet90d] Charles Petzold. A new multiple document interface API simplifies MDI application development. Microsoft Systems Journal, 5(4):53–??, July 1990. CODEN MSJOED. ISSN 0889-9932.

Petzold:1991:MEW

[Pet91a] Charles Petzold. The multimedia extensions for Windows – enhanced sound and video for the PC. Microsoft Systems Journal, 6(2):19–??, March 1991. CODEN MSJOED. ISSN 0889-9932.

Petzold:1991:WTV

[Pet91b] Charles Petzold. Windows version 3.1– hello to TrueType, OLE, and easier DDE; farewell to real mode. Microsoft Systems Journal, 6(5): 17–??, September 1991. CODEN MSJOED. ISSN 0889-9932.

Petzold:1992:CB

[Pet92a] Charles Petzold. The case for 32 bits. Microsoft Systems Journal, 7(4):32–??, July 1992. CODEN MSJOED. ISSN 0889-9932.

Petzold:1992:GCA

[Pet92b] Charles Petzold. GDI comes of age: Exploring the 32-bit graphics of Windows NT. Microsoft Systems Journal, 7(5):41–??, September 1992. CODEN MSJOED. ISSN 0889-9932.

Petzold:1993:VWT

[Pet93] Charles Petzold. Video for Windows brings interleaved audio and full-motion digital video to the PC. Microsoft Systems Journal, 8 (1):41–??, January 1993. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1993:LYC

[Pie93a] Matt Pietrek. Liposuction your corpulent executables and remove excess fat. Microsoft Systems Journal, 8(7):15–??, July 1993. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1993:LUH

[Pie93b] Matt Pietrek. A look under the hood of the Windows 3.1 global heap and the functions that maintain it. Microsoft Systems Journal, 8(3):39–45, 48–52, March 1993. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1993:RPC

[Pie93c] Matt Pietrek. Run privileged code from your Windows-based program using call gates. Microsoft Systems Journal, 8(5):29–??, May 1993. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1993:WDB

[Pie93d] Matt Pietrek. What’s the difference between Windows and tasks in Windows 3.1? Microsoft Systems Journal, 8(10):17–??, October 1993. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1994:PIP

[Pie94a] Matt Pietrek. Peering inside the PE: A tour of the Win32 portable executable file format. Microsoft Systems Journal, 9(3): 15–??, March 1994. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1994:WTQa

[Pie94b] Matt Pietrek. Windows Q and A.

Microsoft Systems Journal, 9(2): 79–??, February 1994. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1994:WTQb

[Pie94c] Matt Pietrek. Windows Q and A. Microsoft Systems Journal, 9 (5):63–??, May 1994. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1995:SBC

[Pie95a] M. Pietrek. Stepping up to 32 bits: Chicago’s process, thread, and memory management. Microsoft Systems Journal, 10(4): 21–??, April 1995. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1995:ETN

[Pie95b] Matt Pietrek. An exclusive tour of the new TOOLHELP32 functions for Windows 95. Microsoft Systems Journal, 10(9):15– ??, September 1995. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1995:UHa

[Pie95c] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 10(9): 73–??, September 1995. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1995:UHb

[Pie95d] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 10 (10):83–??, October 1995. CO-DEN MSJOED. ISSN 0889-9932.

Pietrek:1995:UHc

[Pie95e] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 10 (11):77–??, November 1995. CO-DEN MSJOED. ISSN 0889-9932.

Pietrek:1995:UHd

[Pie95f] Matt Pietrek. Under the hood. Microsoft Systems Journal, 10 (12):75–??, December 1995. CO-DEN MSJOED. ISSN 0889-9932.

Pietrek:1995:WQa

[Pie95g] Matt Pietrek. Windows Q and A. Microsoft Systems Journal, 10 (5):79–??, May 1995. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1995:WQb

[Pie95h] Matt Pietrek. Windows Q and A. Microsoft Systems Journal, 10 (6):81–??, June 1995. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1995:WQc

[Pie95i] Matt Pietrek. Windows Q and A. Microsoft Systems Journal, 10(7):83–??, July 1995. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1995:WQd

[Pie95j] Matt Pietrek. Windows Q and A. Microsoft Systems Journal, 10 (8):91–??, August 1995. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1996:PAU

[Pie96a] Matt Pietrek. Poking around under the hood: A programmer’s view of Windows NT(R) 4.0. Microsoft Systems Journal, 11(8): 18–??, August 1996. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1996:RFD

[Pie96b] Matt Pietrek. Remove fatty deposits from your applications using our 32-bit liposuction tools.

Microsoft Systems Journal, 11 (11):19–??, October 1996. CO-DEN MSJOED. ISSN 0889-9932.

Pietrek:1996:UHa

[Pie96c] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 11(1): 87–??, January 1996. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1996:UHb

[Pie96d] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 11(2): 85–??, February 1996. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1996:UHc

[Pie96e] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 11(3): 77–??, March 1996. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1996:UHd

[Pie96f] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 11 (4):75–??, April 1996. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1996:UHe

[Pie96g] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 11 (5):83–??, May 1996. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1996:UHf

[Pie96h] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 11 (6):79–??, June 1996. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1996:UHg

[Pie96i] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 11 (7):75–??, July 1996. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1996:UHh

[Pie96j] Matt Pietrek. Under the hood. Microsoft Systems Journal, 11(8): 79–??, August 1996. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1996:UHi

[Pie96k] Matt Pietrek. Under the hood. Microsoft Systems Journal, 11(9): 77–??, September 1996. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1996:UHj

[Pie96l] Matt Pietrek. Under the hood. Microsoft Systems Journal, 11 (11):75–??, November 1996. CO-DEN MSJOED. ISSN 0889-9932.

Pietrek:1996:UHk

[Pie96m] Matt Pietrek. Under the hood. Microsoft Systems Journal, 11 (12):67–??, December 1996. CO-DEN MSJOED. ISSN 0889-9932.

Pietrek:1997:CCD

[Pie97a] Matt Pietrek. A crash course on the depths of Win32 structured exception handling. Microsoft Systems Journal, 12(1): 41–??, January 1997. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1997:PPNa

[Pie97b] Matt Pietrek. A programmer’s perspective on new system DLL features in Windows NT 5.0, part I. Microsoft Systems Journal, 12 (11):17–??, November 1997. CO-DEN MSJOED. ISSN 0889-9932.

Pietrek:1997:PPNb

[Pie97c] Matt Pietrek. A programmer’s perspective on new system DLL features in Windows NT 5.0, part II. Microsoft Systems Journal, 12 (12):31–??, December 1997. CO-DEN MSJOED. ISSN 0889-9932.

Pietrek:1997:UHa

[Pie97d] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 12(2): 71–??, February 1997. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1997:UHb

[Pie97e] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 12(3): 79–??, March 1997. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1997:UHc

[Pie97f] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 12 (4):73–??, April 1997. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1997:UHd

[Pie97g] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 12 (5):79–??, May 1997. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1997:UHe

[Pie97h] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 12 (6):83–??, June 1997. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1997:UHf

[Pie97i] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 12(8): 73–??, August 1997. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1997:UHg

[Pie97j] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 12(9):

79–??, September 1997. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1997:UHh

[Pie97k] Matt Pietrek. Under the hood. Microsoft Systems Journal, 12 (11):61–??, November 1997. CO-DEN MSJOED. ISSN 0889-9932.

Pietrek:1997:UHi

[Pie97l] Matt Pietrek. Under the hood. Microsoft Systems Journal, 12 (12):67–??, December 1997. CO-DEN MSJOED. ISSN 0889-9932.

Pietrek:1998:UHa

[Pie98a] Matt Pietrek. Under the hood. Microsoft Systems Journal, 13(1): 77–??, January 1998. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1998:UHb

[Pie98b] Matt Pietrek. Under the hood. Microsoft Systems Journal, 13(3): 77–??, March 1998. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1998:UHc

[Pie98c] Matt Pietrek. Under the hood. Microsoft Systems Journal, 13 (4):75–??, April 1998. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1998:UHd

[Pie98d] Matt Pietrek. Under the hood. Microsoft Systems Journal, 13 (5):71–??, May 1998. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1998:UHe

[Pie98e] Matt Pietrek. Under the hood. Microsoft Systems Journal, 13 (6):85–??, June 1998. CODEN MSJOED. ISSN 0889-9932.

Pietrek:1998:UHf

[Pie98f] Matt Pietrek. Under the hood.

Microsoft Systems Journal, 13 (7):57–??, July 1998. CODEN MSJOED. ISSN 0889-9932.

Platt:1997:FAD

[Pla97a] David Platt. Fashionable app designers agree: The free threading model is what’s hot this fall. Microsoft Systems Journal, 12(8): 33–??, August 1997. CODEN MSJOED. ISSN 0889-9932.

Platt:1997:GAW

[Pla97b] David Platt. Give ActiveX-based Web pages a boost with the apartment threading model. Microsoft Systems Journal, 12(2): 17–??, February 1997. CODEN MSJOED. ISSN 0889-9932.

Prosise:1993:MOA

[Pro93a] Jeff Prosise. MS-DOS 6 offers automatic memory configuration and integrated disk compression. Microsoft Systems Journal, 8(5):13–??, May 1993. CODEN MSJOED. ISSN 0889-9932.

Prosise:1993:MQ

[Pro93b] Jeff Prosise. MS-DOS Q and A. Microsoft Systems Journal, 8 (5):80–??, May 1993. CODEN MSJOED. ISSN 0889-9932.

Prosise:1994:MQ

[Pro94a] Jeff Prosise. MS-DOS Q and A. Microsoft Systems Journal, 9 (5):77–??, May 1994. CODEN MSJOED. ISSN 0889-9932.

Prosise:1994:SAC

[Pro94b] Jeff Prosise. Simplify access to complex multimedia devices with the media control interface. Microsoft Systems Journal, 9(4): 55–??, April 1994. CODEN MSJOED. ISSN 0889-9932.

Prosise:1995:AGW

[Pro95a] J. Prosise. Advanced 3-D graphics for Windows NT 3.5: Introducing the OpenGL interface, part I. Microsoft Systems Journal, 10(5):7– ??, May 1995. CODEN MSJOED. ISSN 0889-9932.

Prosise:1995:MQ

[Pro95b] Jeff Prosise. MS-DOS Q and A. Microsoft Systems Journal, 10 (5):101–??, May 1995. CODEN MSJOED. ISSN 0889-9932.

Prosise:1995:PWMc

[Pro95c] Jeff Prosise. Programming Windows 95 with MFC, palt V: Menus, toolbars, and status bars. Microsoft Systems Journal, 10 (11):41–??, November 1995. CO-DEN MSJOED. ISSN 0889-9932.

Prosise:1995:PWMb

[Pro95d] Jeff Prosise. Programming Windows 95 with MFC, part II: Working with DCs, pens, and brushes. Microsoft Systems Journal, 10 (7):39–??, July 1995. CODEN MSJOED. ISSN 0889-9932.

Prosise:1995:PWMa

[Pro95e] Jeff Prosise. Programming Windows 95 with MFC, part IV: Contending with the keyboard. Microsoft Systems Journal, 10(9):35– ??, September 1995. CODEN MSJOED. ISSN 0889-9932.

Prosise:1995:PWMd

[Pro95f] Jeff Prosise. Programming Windows 95 with MFC, part VI: Dialog boxes, property sheets, and controls. Microsoft Systems Journal, 10(12):53–??, December 1995. CODEN MSJOED. ISSN 0889-9932.

Prosise:1995:WSM

[Pro95g] Jeff Prosise. Wake up and smell the MFC: Using the Visual C++ classes and application framework. Microsoft Systems Journal, 10(6):17–??, June 1995. CODEN MSJOED. ISSN 0889-9932.

Prosise:1996:PWMa

[Pro96a] Jeff Prosise. Programming Windows 95 with MFC, part VII: The Document/View Architecture. Microsoft Systems Journal, 11(2):19–??, February 1996. CO-DEN MSJOED. ISSN 0889-9932.

Prosise:1996:PWMb

[Pro96b] Jeff Prosise. Programming Windows 95 with MFC, part VIII: Printing and print previewing. Microsoft Systems Journal, 11 (4):39–??, April 1996. CODEN MSJOED. ISSN 0889-9932.

Prosise:1996:WCa

[Pro96c] Jeff Prosise. Wicked code. Microsoft Systems Journal, 11(8): 97–??, August 1996. CODEN MSJOED. ISSN 0889-9932.

Prosise:1996:WCb

[Pro96d] Jeff Prosise. Wicked code. Microsoft Systems Journal, 11(12):

81–??, December 1996. CODEN MSJOED. ISSN 0889-9932.

Prosis:1997:WC

[Pro97a] Jeff Prosis. Wicked code. Microsoft Systems Journal, 12(4): 79–??, April 1997. CODEN MSJOED. ISSN 0889-9932.

Prosise:1997:WCa

[Pro97b] Jeff Prosise. Wicked code. Microsoft Systems Journal, 12(2): 87–??, February 1997. CODEN MSJOED. ISSN 0889-9932.

Prosise:1997:WCb

[Pro97c] Jeff Prosise. Wicked code. Microsoft Systems Journal, 12(6):99– ??, June 1997. CODEN MSJOED. ISSN 0889-9932.

Prosise:1997:WCc

[Pro97d] Jeff Prosise. Wicked code. Microsoft Systems Journal, 12(8): 85–??, August 1997. CODEN MSJOED. ISSN 0889-9932.

Patch:1991:DCP

[PS91] Ray Patch and Alok Sinha. Designing the client portion of a LAN Manager client-server application. Microsoft Systems Journal, 6(3):89–??, May 1991. CO-DEN MSJOED. ISSN 0889-9932.

Patch:1992:DNP

[PS92] Ray Patch and Alok Sinha. Developing network-aware programs using Windows 3.1 and Net-BIOS. Microsoft Systems Journal, 7(4):57–??, July 1992. CODEN MSJOED. ISSN 0889-9932.

Quinton:1998:WNB

[Qui98] Michelle Quinton. Windows NT 5.0 brings you new telephony development features with TAPI 3.0. Microsoft Systems Journal, 13 (11):37–??, November 1998. CO-DEN MSJOED. ISSN 0889-9932.

Rauch:1996:TDC

[Rau96a] Stephen Rauch. Talk to any database the COM way using the OLE DB interface. Microsoft Systems Journal, 11(7):19–??, July 1996. CODEN MSJOED. ISSN 0889-9932.

Rauch:1996:UBA

[Rau96b] Stephen Rauch. Unified browsing with Active X extensions brings the Internet to your desktop. Microsoft Systems Journal, 11(9):19– ??, September 1996. CODEN MSJOED. ISSN 0889-9932.

Rauch:1997:MDM

[Rau97] Stephen Rauch. Manage data from myriad sources with the universal data access interfaces. Microsoft Systems Journal, 12(9):17– ??, September 1997. CODEN MSJOED. ISSN 0889-9932.

Richter:1998:FSC

[RC98] Jeffrey Richter and Luis Felipe Cabrera. A file system for the 21ST century: Previewing the Windows NT 5.0 file system. Microsoft Systems Journal, 13(11): 19–??, November 1998. CODEN MSJOED. ISSN 0889-9932.

Richter:1992:DEH

[Ric92a] Jeffrey Richter. Drop everything:

How to make your application accept and source drag-and-drop files. Microsoft Systems Journal, 7(3):19–??, May 1992. CODEN MSJOED. ISSN 0889-9932.

Richter:1992:OCI

[Ric92b] Jeffrey Richter. An overview of Codewright, an industrialstrength editor for Windows. Microsoft Systems Journal, 7(5):61– ??, September 1992. CODEN MSJOED. ISSN 0889-9932.

Richter:1992:SKI

[Ric92c] Jeffrey Richter. Simulating keyboard input between programs requires a (key)stroke of genius. Microsoft Systems Journal, 7(8):43– ??, December 1992. CODEN MSJOED. ISSN 0889-9932.

Richter:1992:SYP

[Ric92d] Jeffrey Richter. Streamline your program’s setup with the version control and decompression DLLs. Microsoft Systems Journal, 7(6): 61–??, October 1992. CODEN MSJOED. ISSN 0889-9932.

Richter:1993:CWT

[Ric93a] Jeffrey Richter. Coordinate Win32 threads using manual-reset and auto-reset events. Microsoft Systems Journal, 8(10):35–??, October 1993. CODEN MSJOED. ISSN 0889-9932.

Richter:1993:CMD

[Ric93b] Jeffrey Richter. Creating, managing, and destroying processes and threads under Windows NT. Microsoft Systems Journal, 8(7):55– ??, July 1993. CODEN MSJOED. ISSN 0889-9932.

Richter:1993:IWT

[Ric93c] Jeffrey Richter. An introduction to Win32 heap and virtual memory management routines. Microsoft Systems Journal, 8(3):15–??, March 1993. CODEN MSJOED. ISSN 0889-9932.

Richter:1993:MFW

[Ric93d] Jeffrey Richter. Memory-mapped files in Windows NT simplify file manipulation and data sharing. Microsoft Systems Journal, 8(4):57–??, April 1993. CODEN MSJOED. ISSN 0889-9932.

Richter:1993:SWT

[Ric93e] Jeffrey Richter. Synchronizing Win 32 threads using critical sections, semaphores, and mutexes. Microsoft Systems Journal, 8(8): 27–??, August 1993. CODEN MSJOED. ISSN 0889-9932.

Richter:1994:LYD

[Ric94a] Jeffrey Richter. Load your 32bit DLL into another process’s address space using INJLIB. Microsoft Systems Journal, 9(5):13– ??, May 1994. CODEN MSJOED. ISSN 0889-9932.

Richter:1994:WQ

[Ric94b] Jeffrey Richter. Win32 Q and A.

Microsoft Systems Journal, 9(2): 69–??, February 1994. CODEN MSJOED. ISSN 0889-9932.

Richter:1995:WRQ

[Ric95a] Jeffery Richter. WIN 32 Q and A.

Microsoft Systems Journal, 10(9): 81–??, September 1995. CODEN MSJOED. ISSN 0889-9932.

Richter:1995:AGM

[Ric95b] Jeffrey Richter. Add growable memory-mapped files to your app with our exclusive GMMF API. Microsoft Systems Journal, 10 (10):17–??, October 1995. CO-DEN MSJOED. ISSN 0889-9932.

Richter:1995:WQa

[Ric95c] Jeffrey Richter. Win 32 Q and A. Microsoft Systems Journal, 10(7):85–??, July 1995. CODEN MSJOED. ISSN 0889-9932.

Richter:1995:WQb

[Ric95d] Jeffrey Richter. Win32 Q and A. Microsoft Systems Journal, 10 (11):85–??, November 1995. CO-DEN MSJOED. ISSN 0889-9932.

Richter:1996:EWS

[Ric96a] Jeffrey Richter. Extend the Windows 95 shell with application desktop toolbars. Microsoft Systems Journal, 11(3):35–??, March 1996. CODEN MSJOED. ISSN 0889-9932.

Richter:1996:FYA

[Ric96b] Jeffrey Richter. Fusing your applications to the system through the Windows 95 shell. Microsoft Systems Journal, 11(4):61–??, April 1996. CODEN MSJOED. ISSN 0889-9932.

Richter:1996:WYN

[Ric96c] Jeffrey Richter. What you need to know about “Pegasus,” Microsoft’s upcoming handheld device. Microsoft Systems Journal, 11(11):41–??, October 1996. CO-DEN MSJOED. ISSN 0889-9932.

Richter:1996:W

[Ric96d] Jeffrey Richter. Win 32 Q and A. Microsoft Systems Journal, 11 (11):85–??, November 1996. CO-DEN MSJOED. ISSN 0889-9932.

Richter:1996:WQb

[Ric96e] Jeffrey Richter. Win 32 Q+A.

Microsoft Systems Journal, 11(3): 87–??, March 1996. CODEN MSJOED. ISSN 0889-9932.

Richter:1996:WQd

[Ric96f] Jeffrey Richter. Win 32(R) Q and A. Microsoft Systems Journal, 11(7):83–??, July 1996. CODEN MSJOED. ISSN 0889-9932.

Richter:1996:WQa

[Ric96g] Jeffrey Richter. Win32 Q and A.

Microsoft Systems Journal, 11(1): 93–??, January 1996. CODEN MSJOED. ISSN 0889-9932.

Richter:1996:WQe

[Ric96h] Jeffrey Richter. Win32 Q and A.

Microsoft Systems Journal, 11(9): 91–??, September 1996. CODEN MSJOED. ISSN 0889-9932.

Richter:1996:WQc

[Ric96i] Jeffrey Richter. Win32(R) Q and A. Microsoft Systems Journal, 11 (5):91–??, May 1996. CODEN MSJOED. ISSN 0889-9932.

Richter:1997:DWN

[Ric97a] Jeffrey Richter. Design a Windows NT service to exploit special operating system facilities. Microsoft Systems Journal, 12(10): 19–??, October 1997. CODEN MSJOED. ISSN 0889-9932.

Richter:1997:Wb

[Ric97b] Jeffrey Richter. Win 32. Microsoft Systems Journal, 12(11): 67–??, November 1997. CODEN MSJOED. ISSN 0889-9932.

Richter:1997:WQa

[Ric97c] Jeffrey Richter. Win 32 Q and A. Microsoft Systems Journal, 12 (3):85–??, March 1997. CODEN MSJOED. ISSN 0889-9932.

Richter:1997:WQb

[Ric97d] Jeffrey Richter. Win 32 Q and A. Microsoft Systems Journal, 12 (5):91–??, May 1997. CODEN MSJOED. ISSN 0889-9932.

Richter:1997:Wa

[Ric97e] Jeffrey Richter. Win32. Microsoft Systems Journal, 12(9):91– ??, September 1997. CODEN MSJOED. ISSN 0889-9932.

Richter:1998:CPM

[Ric98a] Jeffrey Richter. Custom performance monitoring for your Windows NT applications. Microsoft Systems Journal, 13(8):17–??, August 1998. CODEN MSJOED. ISSN 0889-9932.

Richter:1998:MWN

[Ric98b] Jeffrey Richter. Manipulate Windows NT services by writing a service control program. Microsoft Systems Journal, 13(2): 55–??, February 1998. CODEN MSJOED. ISSN 0889-9932.

Richter:1998:Wa

[Ric98c] Jeffrey Richter. Win 32. Microsoft Systems Journal, 13(1):

85–??, January 1998. CODEN MSJOED. ISSN 0889-9932.

Richter:1998:Wb

[Ric98d] Jeffrey Richter. Win32. Microsoft Systems Journal, 13(3): 83–??, March 1998. CODEN MSJOED. ISSN 0889-9932.

Richter:1998:Wc

[Ric98e] Jeffry Richter. Win32. Microsoft Systems Journal, 13(6):95– ??, June 1998. CODEN MSJOED. ISSN 0889-9932.

Robbins:1997:B

[Rob97a] John Robbins. Bugslayer. Microsoft Systems Journal, 12(12): 89–??, December 1997. CODEN MSJOED. ISSN 0889-9932.

Robbins:1997:IWI

[Rob97b] John Robbins. Ingenious ways to implement multiple threads in Visual Basic 5.0, part I. Microsoft Systems Journal, 12(8):17–??, August 1997. CODEN MSJOED. ISSN 0889-9932.

Robbins:1997:IBA

[Rob97c] John Robbins. Introducing the bugslayer: Annihilating bugs in an application near you. Microsoft Systems Journal, 12(10): 35–??, October 1997. CODEN MSJOED. ISSN 0889-9932.

Robbins:1997:MTV

[Rob97d] John Robbins. Multiple threads in visual basic 5.0, part II: Writing a Win32 debugger. Microsoft Systems Journal, 12(9):45– ??, September 1997. CODEN MSJOED. ISSN 0889-9932.

Robbins:1998:Ba

[Rob98a] John Robbins. Bugslayer. Microsoft Systems Journal, 13(4): 89–??, April 1998. CODEN MSJOED. ISSN 0889-9932.

Robbins:1998:Bb

[Rob98b] John Robbins. Bugslayer. Microsoft Systems Journal, 13(6):99– ??, June 1998. CODEN MSJOED. ISSN 0889-9932.

Roberts:1998:KEY

[Rob98c] Scot Roberts. Keeping an eye on your browser by monitoring Internet Explorer 4.0 events. Microsoft Systems Journal, 13(6):69– ??, June 1998. CODEN MSJOED. ISSN 0889-9932.

Rozak:1996:TYC

[Roz96] Mike Rozak. Talk to your computer and have it answer back with the Microsoft speech API. Microsoft Systems Journal, 11(1): 19–??, January 1996. CODEN MSJOED. ISSN 0889-9932.

Reed:1997:MTS

[RTT97] Dave Reed, Tracey Trewin, and Mai-Ian Tomsen. Microsoft Transaction Server helps you write scalable, distributed Internet apps. Microsoft Systems Journal, 12(8): 51–??, August 1997. CODEN MSJOED. ISSN 0889-9932.

Smetak:1997:DRO

[SC97] Ed Smetak and Jean Caputo. Dynamic runtime objects: Building applications your users can modify at runtime. Microsoft Systems Journal, 12(7):49–??, July 1997.

CODEN MSJOED. ISSN 0889-9932.

Schulman:1992:EDV

[Sch92a] Andrew Schulman. Exploring demand-paged virtual memory in Windows enhanced mode. Microsoft Systems Journal, 7(8):17– ??, December 1992. CODEN MSJOED. ISSN 0889-9932.

Schulman:1992:GAD

[Sch92b] Andrew Schulman. Go anywhere and do anything with 32-bit virtual device drivers for Windows. Microsoft Systems Journal, 7(6): 15–??, October 1992. CODEN MSJOED. ISSN 0889-9932.

Schulman:1993:ALB

[Sch93a] Andrew Schulman. At last-write bona fide 32-bit programs that run on Windows 3.1 using Win32s. Microsoft Systems Journal, 8(4): 15–??, April 1993. CODEN MSJOED. ISSN 0889-9932.

Schulman:1993:CVF

[Sch93b] Andrew Schulman. Call VxD functions and VMM services easily using our generic VxD. Microsoft Systems Journal, 8(2):17– ??, February 1993. CODEN MSJOED. ISSN 0889-9932.

Schmid:1997:IMD

[Sch97] Robert Schmid. Investigating multilanguage debugging and the new IDEs. Microsoft Systems Journal, 12(5):51–??, May 1997. CODEN MSJOED. ISSN 0889-9932.

Sinha:1997:BSA

[SCP97] Alok Sinha, Don Clore, and Dale Phurrough. Behind the scenes at MSN 2.0: Architecting an Internet-based online service. Microsoft Systems Journal, 12(4): 31–??, April 1997. CODEN MSJOED. ISSN 0889-9932.

Shaw:1990:APM

[Sha90a] Richard Hale Shaw. Accessing Presentation Manager facilities from within OS/2 kernel applications. Microsoft Systems Journal, 5(1):01–??, January 1990. CO-DEN MSJOED. ISSN 0889-9932.

Shaw:1990:BPC

[Sha90b] Richard Hale Shaw. Based pointers: Combining far pointer addressability and the small size of near pointers. Microsoft Systems Journal, 5(5):51–??, September 1990. CODEN MSJOED. ISSN 0889-9932.

Shaw:1990:IDP

[Sha90c] Richard Hale Shaw. Implementing DDE with Presentation Manager object Windows. Microsoft Systems Journal, 5(3):81–??, May 1990. CODEN MSJOED. ISSN 0889-9932.

Shaw:1990:PDF

[Sha90d] Richard Hale Shaw. Providing DDE facilities to OS/2 kernel applications. Microsoft Systems Journal, 5(4):65–??, July 1990. CODEN MSJOED. ISSN 0889-

Shaw:1990:TID

[Sha90e] Richard Hale Shaw. ToolBook, and Interactive Development Environment for Windows. Microsoft Systems Journal, 5(5):93– ??, September 1990. CODEN MSJOED. ISSN 0889-9932.

Shaw:1991:AOT

[Sha91a] Richard Hale Shaw. Analyzing the optimization techniques compilers use to transform your C code. Microsoft Systems Journal, 6(2):29–??, March 1991. CODEN MSJOED. ISSN 0889-9932.

Shaw:1991:BCP

[Sha91b] Richard Hale Shaw. Borland C++ 2.0 provides a comprehensive Windows development environment. Microsoft Systems Journal, 6(3):45–??, May 1991. CO-DEN MSJOED. ISSN 0889-9932.

Shaw:1991:QWT

[Sha91c] Richard Hale Shaw. QuickC for Windows, a complete graphical environment for easier Windows programming. Microsoft Systems Journal, 6(6):65–??, November 1991. CODEN MSJOED. ISSN 0889-9932.

Shaw:1992:DML

[Sha92a] Richard Hale Shaw. Do more with less code: Pare down your Windows-based applications using C++. Microsoft Systems Journal, 7(5):15–??, September 1992. CO-DEN MSJOED. ISSN 0889-9932.

Shaw:1992:PMC

[Sha92b] Richard Hale Shaw. A preview of Microsoft C/C++ 7 and the Microsoft Foundation Classes for Windows. Microsoft Systems Journal, 7(2):43–??, March 1992. CODEN MSJOED. ISSN 0889-9932.

Shaw:1992:PWT

[Sha92c] Richard Hale Shaw. Programming Windows with a touch of class: The Microsoft Foundation Classes. Microsoft Systems Journal, 7(7):19–??, November 1992. CODEN MSJOED. ISSN 0889-9932.

Shaw:1993:SNC

[Sha93] Richard Hale Shaw. Surveying new capabilities of the Microsoft Foundation Class library version 2.0. Microsoft Systems Journal, 8(6):41–??, June 1993. CODEN MSJOED. ISSN 0889-9932.

Sherlock:1991:PAW

[She91] Thomas Sherlock. Porting 32bit applications to Windows 3.0 with the WINMEM32 library. Microsoft Systems Journal, 6(2): 57–??, March 1991. CODEN MSJOED. ISSN 0889-9932.

Shepherd:1997:VCS

[She97] George Shepherd. Visual C++5.0 simplifies the process for developing and using COM objects. Microsoft Systems Journal, 12(5):37– ??, May 1997. CODEN MSJOED. ISSN 0889-9932.

Skonnard:1998:HDR

[Sko98] Aaron Skonnard. How to design reusable HTTP components by exploiting Winlet and COM. Microsoft Systems Journal, 13(6):45– ??, June 1998. CODEN MSJOED. ISSN 0889-9932.

Smith:1991:EDD

[Smi91] Gordon Smith. Embedded device drivers simplify the support of unusual devices under Windows. Microsoft Systems Journal, 6(3):63– ??, May 1991. CODEN MSJOED. ISSN 0889-9932.

Smith:1992:SCA

[Smi92] Gordon S. Smith. Simplify clientserver applications using the DDE management library. Microsoft Systems Journal, 7(8):61–??, December 1992. CODEN MSJOED. ISSN 0889-9932.

Shepherd:1997:VPa

[SW97a] George Shepherd and Scot Wingo.

Visual programmer. Microsoft Systems Journal, 12(4):87–??, April 1997. CODEN MSJOED. ISSN 0889-9932. URL http://

www.microsoft.com/msj/defaultframe.
asp?page=/msj/0497/visualprog/
visualprog0497.htm&nav=/msj/
0497/newnav.htm.

Shepherd:1997:VPb

[SW97b] George Shepherd and Scot Wingo.

The visual programmer. Microsoft Systems Journal, 12(6): 105–??, June 1997. CODEN MSJOED. ISSN 0889-9932.

Shepherd:1997:VPc

[SW97c] George Shepherd and Scot Wingo.

The visual programmer. Microsoft Systems Journal, 12(8): 101–??, August 1997. CODEN MSJOED. ISSN 0889-9932.

Shepherd:1998:VPa

[SW98a] Shepherd and Wing. Visual programmer. Microsoft Systems Journal, 13(4):83–??, April 1998. CODEN MSJOED. ISSN 0889-9932.

Shepherd:1998:VPb

[SW98b] Shepherd and Wingo. Visual programmer. Microsoft Systems Journal, 13(7):67–??, July 1998. CODEN MSJOED. ISSN 0889-9932.

Thielen:1991:LFT

[Thi91a] David Thielen. Less filling, tastes great: A programmer’s survey of MS-DOS version 5. Microsoft Systems Journal, 6(4):17–??, July 1991. CODEN MSJOED. ISSN 0889-9932.

Thielen:1991:UTM

[Thi91b] David Thielen. Unconventional thoughts on managing PC software development. Microsoft Systems Journal, 6(3):111–??, May 1991. CODEN MSJOED. ISSN 0889-9932.

Thielen:1992:SFW

[Thi92] David Thielen. Safer functions for working with MS-DOS files. Microsoft Systems Journal, 7(4):79– ??, July 1992. CODEN MSJOED. ISSN 0889-9932.

Thielen:1994:PPV

[Thi94] David Thielen. A programmer’s potpourri of value-added tools for Visual C++ and MFC. Microsoft Systems Journal, 9(3): 71–??, March 1994. CODEN MSJOED. ISSN 0889-9932.

Trupin:1993:DSH

[Tru93a] Joshua Trupin. Dr. Strangerelease; or how I learned to stop worrying and love version control systems. Microsoft Systems Journal, 8(5):43–??, May 1993. CODEN MSJOED. ISSN 0889-9932.

Trupin:1993:HDA

[Tru93b] Joshua Trupin. Harnessing the database access and OLE support capabilities of Visual Basic 3.0. Microsoft Systems Journal, 8(9): 15–??, September 1993. CODEN MSJOED. ISSN 0889-9932.

Trupin:1993:IJP

[Tru93c] Joshua Trupin. It’s not just for prototypes any more: Extra power and controls enhance Visual Basic 2.0. Microsoft Systems Journal, 8 (1):15–??, January 1993. CODEN MSJOED. ISSN 0889-9932.

Trupin:1994:AIV

[Tru94] Joshua Trupin. Application interoperability with Visual Basic for applications and OLE 2.0. Microsoft Systems Journal, 9(2): 29–??, February 1994. CODEN MSJOED. ISSN 0889-9932.

Trupin:1995:BRD

[Tru95a] Joshua Trupin. Borland Delphi:

A new entry in the pantheon of visual programming tools. Microsoft Systems Journal, 10(6):37– ??, June 1995. CODEN MSJOED. ISSN 0889-9932.

Trupin:1995:VBP

[Tru95b] Joshua Trupin. Visual Basic 4.0 provides easy entry to the advanced features of Windows(9)

95. Microsoft Systems Journal, 10(10):35–??, October 1995. CO-DEN MSJOED. ISSN 0889-9932.

Trupin:1996:VPa

[Tru96a] Joshua Trupin. The visual programmer. Microsoft Systems Journal, 11(1):81–??, January 1996. CODEN MSJOED. ISSN 0889-9932.

Trupin:1996:VPb

[Tru96b] Joshua Trupin. The visual programmer. Microsoft Systems Journal, 11(2):93–??, February 1996. CODEN MSJOED. ISSN 0889-9932.

Trupin:1996:VPc

[Tru96c] Joshua Trupin. The visual programmer. Microsoft Systems Journal, 11(4):103–??, April 1996. CODEN MSJOED. ISSN 0889-9932.

Trupin:1996:VPd

[Tru96d] Joshua Trupin. The visual programmer. Microsoft Systems Journal, 11(5):95–??, May 1996. CODEN MSJOED. ISSN 0889-9932.

Trupin:1996:VPF

[Tru96e] Joshua Trupin. The visual programmer fingers WinSock functions from Visual Basic. Microsoft Systems Journal, 11(3): 19–??, March 1996. CODEN MSJOED. ISSN 0889-9932.

Trupin:1996:VPP

[Tru96f] Joshua Trupin. The visual programmer puts ActiveX(TM) document objects through their paces. Microsoft Systems Journal, 11(6):55–??, June 1996. CODEN MSJOED. ISSN 0889-9932.

Trupuin:1996:VP

[Tru96g] Joshua Trupuin. The visual programmer. Microsoft Systems Journal, 11(12):75–??, December 1996. CODEN MSJOED. ISSN 0889-9932.

Umeno:1998:TAP

[Ume98] Hiroo Umeno. For the telephony API, press 1; for unimodem, press 2; or stay on the line. Microsoft Systems Journal, 13(4):33–??, April 1998. CODEN MSJOED. ISSN 0889-9932.

Weisz:1996:FAT

[Wei96] Russell Weisz. First aid for the thread-impaired: Using multiple threads with MFC. Microsoft Systems Journal, 11(12):19–??, December 1996. CODEN MSJOED. ISSN 0889-9932.

Weisz:1997:MFA

[Wei97] Russell Weisz. More first aid for the thread impaired: Cool ways to take advantage of multithreading. Microsoft Systems Journal, 12(7):33–??, July 1997. CODEN MSJOED. ISSN 0889-9932.

Welch:1990:BEC

[Wel90a] Kevin P. Welch. Building an extensible clipboard viewer through the use of dynamic-link libraries. Microsoft Systems Journal, 5(2): 27–??, March 1990. CODEN MSJOED. ISSN 0889-9932.

Welch:1990:CDR

[Wel90b] Kevin P. Welch. Creating a data-entry and reporting library for Windows 3.0: Part II. Microsoft Systems Journal, 5(6):99– ??, November 1990. CODEN MSJOED. ISSN 0889-9932.

Welch:1990:CMD

[Wel90c] Kevin P. Welch. Creating a multipage data-entry and reporting library for Windows version 3.0. Microsoft Systems Journal, 5(5): 15–??, September 1990. CODEN MSJOED. ISSN 0889-9932.

Welch:1990:EWT

[Wel90d] Kevin P. Welch. Extending the Windows 3.0 interface with installable custom controls. Microsoft Systems Journal, 5(4):29–??, July 1990. CODEN MSJOED. ISSN 0889-9932.

Welch:1990:TTO

[Wel90e] Kevin P. Welch. Techniques for tuning and optimization of PM and Windows applications. Microsoft Systems Journal, 5(1): 47–??, January 1990. CODEN MSJOED. ISSN 0889-9932.

Welch:1990:UOM

[Wel90f] Kevin P. Welch. Using objectoriented methodologies in Windows applications. Microsoft Systems Journal, 5(3):63–??, May 1990. CODEN MSJOED. ISSN 0889-9932.

Welch:1991:WTW

[Wel91] Kevin P. Welch. What’s in there? Windows 3.0 executable and resource file formats revealed! Microsoft Systems Journal, 6(5):43– ??, September 1991. CODEN MSJOED. ISSN 0889-9932.

Welch:1992:CGI

[Wel92] Kevin P. Welch. Creating graphics import filter libraries for Windows applications. Microsoft Systems Journal, 7(1):101–??, January 1992. CODEN MSJOED. ISSN 0889-9932.

Williams:1993:MIM

[Wil93] Sara Williams. An MSJ interview with Microsoft’s chief architect of OLE, Tony Williams. Microsoft Systems Journal, 8(10):55–??, October 1993. CODEN MSJOED. ISSN 0889-9932.

Wood:1990:CWT

[Woo90] Robert A. Wood. Creating Windows list boxes that support virtualized scrolling. Microsoft Systems Journal, 5(3):47–??, May 1990. CODEN MSJOED. ISSN 0889-9932.

Yao:1990:WTM

[Yao90] Paul Yao. Windows 3.0 memory management: Supporting disparate 80x86 architectures. Microsoft Systems Journal, 5(6):35– ??, November 1990. CODEN MSJOED. ISSN 0889-9932.

Yao:1991:CWT

[Yao91a] Paul Yao. Careful Windows resource allocation and cleanup improves application hygiene. Microsoft Systems Journal, 6(5):31– ??, September 1991. CODEN MSJOED. ISSN 0889-9932.

Yao:1991:IWT

[Yao91b] Paul Yao. Improve Windows application memory use with subsegment allocation and custom resources. Microsoft Systems Journal, 6(1):75–??, January 1991. CODEN MSJOED. ISSN 0889-9932.

Yao:1991:TPW

[Yao91c] Paul Yao. Tuning the performance of Windows and OS/2 programs with MicroQuill’s segmentor. Microsoft Systems Journal, 6 (2):49–??, March 1991. CODEN MSJOED. ISSN 0889-9932.

Yao:1991:WTA

[Yao91d] Paul Yao. Windows 32-Bit API gives developers advanced operating system capabilities. Microsoft Systems Journal, 6(6):15– ??, November 1991. CODEN MSJOED. ISSN 0889-9932.

Yao:1992:EPU

[Yao92a] Paul Yao. Explore previously uncharted areas of Windows using the ToolHelp library. Microsoft Systems Journal, 7(3):37–??, May 1992. CODEN MSJOED. ISSN 0889-9932.

Yao:1992:IWN

[Yao92b] Paul Yao. An introduction to Windows NT memory management fundamentals. Microsoft Systems Journal, 7(4):41–??, July 1992. CODEN MSJOED. ISSN 0889-9932.

Yao:1998:MWC

[Yao98] Paul Yao. Microsoft Windows CE 2.0: It’s not just for handheld PCs any more. Microsoft Systems Journal, 13(5):19–??, May 1998. CODEN MSJOED. ISSN 0889-9932.

Zimmerman:1997:CAE

[Zim97a] Steve Zimmerman. Create apps that are easily extensible with our smart “snap- ins” methodology. Microsoft Systems Journal, 12(7):17–??, July 1997. CODEN MSJOED. ISSN 0889-9932.

Zimmerman:1997:EDS

[Zim97b] Steve Zimmerman. Extend Developer Studio 97 with your own add-ins, macros, and wizards. Microsoft Systems Journal, 12(9):29– ??, September 1997. CODEN MSJOED. ISSN 0889-9932.